### v1

# Censored extreme value estimation

Abstract
A novel and comprehensive methodology designed to tackle the challenges posed by extreme values in the context of random censorship is introduced. The main focus is the analysis of integrals based on the product-limit estimator of normalized top-order statistics, denoted extreme Kaplan–Meier integrals. These integrals allow for transparent derivation of various important asymptotic distributional properties, offering an alternative approach to conventional plug-in estimation methods. Notably, this methodology demonstrates robustness and wide applicability within the scope of max-domains of attraction. An additional noteworthy by-product is the extension of residual estimation of extremes to encompass all max-domains of attraction, which is of independent interest.
62G32,
62N02,
62E20,
Censored extremes,
Kaplan-Meier estimator,
Random censoring,
Extreme value index,
Residual estimators,
Regular variation,
Tail inference,
keywords:
[class=MSC]
keywords:
[…]
and
1 Introduction
The foundational statistical work in fully-observed extreme value theory is by now classic and well-understood in the univariate case, see [2, 19, 15], and to a large degree for univariate time series, as can be found in [30]. Similarly, the pillars of survival analysis have been laid out since the seventies, in different flavors, with the counting process point of view being one of the the most popular approaches, and with martingale techniques playing an ubiquitous role, see [24, 1].
A staple of nonparametric estimation of the survival function for randomly censored time-to-event data is the Kaplan–Meier product-limit estimator [29], which has received continued attention through the decades. In [13] the rate of strong uniform consistency for the product-limit estimator was studied, and in [27] a law of the iterated logarithm was provided. The functional point of view has been explored in [16], providing functional limit laws for the increment functions of empirical processes based on randomly right-censored data. Concentrating on the tail empirical process of a fully-observed sample, the results of [10] provide the main tools to assert asymptotic functional distributional properties for extreme estimators. The empirical process approach has also resulted in fruitful limit theorems for time-series, as treated in [30].
Integral functions of the empirical distribution function were studied in [11] in terms of normal and stable limits. The generalization of this idea turned out to be paramount for survival analysis, integrating with respect to the Kaplan–Meier estimator. Indicator functions were first treated in [36] and [41], while the mean residual lifetime, corresponding to integrating the identity function, was treated in [40]. In the latter, the main mathematical technique was integration by parts. In [39] arbitrary integrals with respect to the Kaplan–Meier estimator were considered, establishing strong convergence through martingale techniques, and then [38] provided the central limit theorem using U-statistics. These results provide a general framework for dealing with distributional estimators from censored samples, and importantly, seamlessly deal with convergence on the largest possible support – the maximal interval – by appealing to U-statistics theory. Convergence of the product-limit estimator on the maximal interval was first dealt with successfully in [41], see also [42] for random truncation.
Integral functions with respect to the empirical distribution function of top order statistics has also received interest. Such integrals can often be cast as sums of extreme values from regularly varying distributions, which was studied in [14] for regularly-varying tails and in [12] for arbitrary tail behavior. Upon replacement of the empirical distribution function with the Kaplan–Meier estimator of top-order statistics of a censored sample, the analysis of integral functions is no longer transparent with empirical processes theory, nor using the sums approach. This paper is devoted to studying precisely these integral functions by using the elegant U-statistics approach of [38], together with machinery from extreme value theory.
More precisely, this paper proposes the analysis of randomly right-censored extremes through Kaplan–Meier integrals of normalized top-order statistics. In order to generalize the approach of [38] to the extreme case, the theory of U-statistics [35] and of residual estimation of [34] are of utmost importance. However, they are not directly applicable, the former having to be adapted to accommodate upper order statistics (see [32] for a related concept), and the latter requiring extension to the Gumbel and Weibull max-domains of attraction. The result provides a unifying and transparent treatment of extreme estimators and their convergence, imposing natural conditions. Note that extreme value index estimators derived from extreme Kaplan–Meier integrals have similar form and in some cases equal asymptotic variance as the ones treated in [6], though the construction principle is rather different and our theory allows for the construction of further estimators. In particular we provide a detailed account on a novel construction of the censored version of the moment estimator [17] of the extreme value index for all max-domains of attraction.
The main theoretical contributions of the paper can be summarized as follows:
  1. 1.
    Suitable conditions are established under which extreme Kaplan–Meier integrals are weakly consistent and asymptotically normal within the Fréchet max-domain of attraction, that is for distributions with regularly varying tails. Asymptotic normality is provided in several steps, with only the last one requiring a second order condition. The main unavoidable restriction is a second (plus epsilon) moment condition for the limit variance to exist, and in particular there are no requirements on the rate of uniform convergence of the asymptotic censoring proportion.
  2. 2.
    Residual estimation theory is then generalized to all max-domains of attraction, and in particular it is found that convergence rates vary across domains. There results are presented for fully-observed data and are of independent interest. The exposition is in terms of so-called candidate functions, which themselves must satisfy a second-order type condition.
  3. 3.
    The results on residual estimators is then used to extend the distributional convergence of extreme Kaplan–Meier-type integrals for arbitrary max-domains of attraction. Here, a regularly varying index of certain cdf’s plays a central role. Such index can be identified as the asymptotic non-censoring proportion.
  4. 4.
    Finally, we provide full asymptotics for a novel censored version of the moment estimator of [17], using all the main components of our theory. Some assumptions from the general theory are relaxable for this application and thus some more technical proofs are provided for precisely this purpose.
The cross-fertilization of extreme value theory and survival analysis has been explored before in various contexts. We provide an account of some of the most important contributions. In [3] the Pareto index estimation was considered for the first time under random right-censorship framework (see also [4, 18] covering all max-domains of attraction), which has since produced some generalizations, e.g. using covariates in [31], truncation in [25], trimming in [9], and bias-reduction in [5, 26]. Another approach, which neatly formulates the censoring problem with extremes as a path-wise problem, was proposed in [43] through constructions based on expected values. Subsequently, a partial result on the asymptotic distribution was recently provided in [6]. However, their method of proof is not generalizable to our setting. In terms of missing extreme data, the main reference is [44], and finally the machinery of extreme value theory has recently been imported to survival analysis in a different context, namely for cure-rate modelling in the case of insufficient follow-up, see [20, 21, 22].
Organisation of the paper. In Section 2 we introduce notation and results from extreme value theory and survival analysis relevant to our setup. Section 3 is devoted to the study of extreme Kaplan–Meier integrals when both event and censoring variables have distributions with regularly varying tails; Theorem 3.3 provides a general decomposition, which is exploited in Theorem 3.4 for consistency and in a series of asymptotic normality results culminating in Corollary 3.10. Residual estimators and their applications are then studied in Section 4 with no max-domain constraints and extensions of the results of the previous section are provided; we highlight Theorem 4.11 and Theorem 4.12 as our main extensions. Finally, Section 5 provides a prime application of the previous sections, analysing in detail an extreme value index estimator for censored data and all max-domains of attraction; we provide in Theorem 5.1 consistency and in Theorem 5.2 asymptotic normality. Proofs may be found in Appendices A, B and C.
2 Preliminaries and setting
2.1 Random censoring
Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with common cdf $F.$
Let $\varphi:\amsmathbb{R}\to\amsmathbb{R}$ be some measurable function which is $F$-square integrable, that is $\int\varphi^{2}\mathop{}\!\mathrm{d}F<\infty$. Then, an application of the central limit theorem (CLT) for $\{\varphi(X_{i})\}_{i=1}^{n}$ can be written as
$\sqrt{n}\int\varphi\,\mathop{}\!\mathrm{d}(F_{n}-F)\stackrel{{\scriptstyle d}}{{\to}}N(0,\sigma_{\varphi}^{2}),$ (2.1)
where $F_{n}$ is the empirical cdf of $\{X_{i}\}_{i=1}^{n}$ and $\sigma_{\varphi}^{2}=\int\varphi^{2}\mathop{}\!\mathrm{d}F-\left(\int\varphi\,\mathop{}\!\mathrm{d}F\right)^{2}.$ We now consider a random right-censoring assumption, that is, we assume that we only can observe
$Z_{i}=\min(X_{i},Y_{i})\>\>\>\text{ and }\>\>\>\delta_{i}=I(X_{i}\leq Y_{i}),$
where $Y_{1},\ldots,Y_{n}$ is an i.i.d. censoring sequence with common cdf $G$ independent of $\{X_{i}\}_{i=1}^{n}$. Notice that $\delta_{i}$ indicates whether $X_{i}$ has been observed or not. Let $H$ denote the cdf of $Z_{1}$, which by independence of $X_{1}$ and $Y_{1},$ is given by
$1-H=(1-F)(1-G).$ (2.2)
A consistent estimator of $F$, [41], is now provided under by the Kaplan–Meier product-limit estimator $\amsmathbb{F}_{n}$ instead of $F_{n}$ in (2.1), defined by
$1-\amsmathbb{F}_{n}(x)=\prod_{i=1}^{n}\left[1-\frac{\delta_{[i:n]}}{n-i+1}\right]^{I(Z_{i,n}\leq x)}.$ (2.3)
Here $Z_{1,n}\leq\ldots\leq Z_{n,n}$ are the order statistics of $\{Z_{i}\}$ and $\{\delta_{[i:n]}\}_{i=1}^{n}$ are the concomitants of $\{Z_{i,n}\}_{i=1}^{n},$ that is $\delta_{[i:n]}=\delta_{j}$ if $Z_{i,n}=Z_{j}.$
More precisely, write $F\{a\}=F(a)-F(a-)$, such that $F\{a\}>0$ iff $a$ is an atom of $F$. Let $A$ be the set of all atoms of $H,$ possibly empty. Under $\int|\varphi|\mathop{}\!\mathrm{d}F<\infty,$ Stute and Wang [39] showed that
$\int\varphi\,\mathop{}\!\mathrm{d}\amsmathbb{F}_{n}\stackrel{{\scriptstyle a.s.}}{{\to}}\int\varphi\,\mathop{}\!\mathrm{d}\widetilde{F},\quad n\to\infty,$
where $\tau_{H}=\inf\{x:H(x)=1\}$ and
$\widetilde{F}(x)=\left\{\begin{array}[]{cc}F(x),&\text{ if }x<\tau_{H};\\ F(\tau_{H}-)+I(\tau_{H}\in A)F\{\tau_{H}\},&\text{ otherwise.}\end{array}\right.$ (2.4)
Theorem 1.1 in [38] then states that under some mild regularity conditions imposed on $\varphi,$ $F$ and $G$, the Kaplan–Meier integral $\int\varphi\,\mathop{}\!\mathrm{d}\amsmathbb{F}_{n}$ can be represented as a sum of i.i.d. random variables and a summand $R_{n}$ vanishing as $o_{P}(n^{-1/2})$.
Introduce the subdistribution functions, for $z\in\amsmathbb{R}$,
$H^{j}(z)$ $={\amsmathbb{P}}(Z\leq z,\delta=j),\quad j=0,1,$
where $(Z,\delta)\stackrel{{\scriptstyle d}}{{=}}(Z_{1},\delta_{1}).$
The following three functions are necessary to state a key asymptotic decomposition of $\int\varphi\,\mathop{}\!\mathrm{d}\amsmathbb{F}_{n}$, and are stated here for future reference.
$\gamma_{0}(x)=\exp\left\{\int_{z<x}\frac{H^{0}(\mathop{}\!\mathrm{d}z)}{1-H(z)}\right\},$ (2.5)
$\gamma_{1}(x)=\frac{1}{1-H(x)}\int_{x<z}\varphi(z)\gamma_{0}(z)H^{1}(\mathop{}\!\mathrm{d}z),$ (2.6)
and
$\gamma_{2}(x)=\int\int\frac{I(v<x,v<z)\varphi(z)\gamma_{0}(z)}{(1-H(v))^{2}}H^{0}(\mathop{}\!\mathrm{d}v)H^{1}(\mathop{}\!\mathrm{d}z).$ (2.7)
These functions can be simplified if $F$ and $G$ are continuous, for instance
$\gamma_{0}(x)=\frac{1}{1-G(x)},\quad x<\tau_{H}.$ (2.8)
The decomposition in turn implies the central limit theorem under random censorship (see Corollary 1.2, [38]),
$\sqrt{n}\int\varphi\,\mathop{}\!\mathrm{d}(\amsmathbb{F}_{n}-\widetilde{F})\stackrel{{\scriptstyle d}}{{\to}}N(0,\sigma_{\varphi}^{2}),$ (2.9)
where now $\sigma_{\varphi}^{2}={\rm var}\{\varphi(Z)\gamma_{0}(Z)\delta+\gamma_{1}(Z)(1-\delta)-\gamma_{2}(Z)\}.$
2.2 Extreme value theory
We are interested in analyzing the extreme value behaviour of the datapoints $X_{1},\dots,X_{n}$, which is possible through the study of $Z_{i,n}$ and $\delta_{[i:n]}$.
Define for any cdf $F$ its survival function as $\bar{F},$ that is $\bar{F}=1-F,$ and its tail counterpart as
$F^{t}(x)=\frac{F(xt)-F(t)}{1-F(t)},\quad x\geq 1,$ (2.10)
which is itself a distribution function. Further, let $\gamma^{t}_{0},$ $\gamma^{t}_{1}$ and $\gamma^{t}_{2}$ are defined similarly to (2.5), (2.6) and (2.7), by replacing $H,H^{0}$ and $H^{1}$ with $H^{t},$ $H^{0,t}$ and $H^{1,t},$ respectively, where
$H^{0,t}(x)=\int_{1}^{x}(1-F^{t}(z))G^{t}(\mathop{}\!\mathrm{d}z)\>\>\>\text{ and }\>\>\>H^{1,t}(x)=\int_{1}^{x}(1-G^{t}(z-))F^{t}(\mathop{}\!\mathrm{d}z).$ (2.11)
Finally, for future reference, define
$C^{t}(x)=\int_{1}^{\infty}\frac{G^{t}(\mathop{}\!\mathrm{d}y)}{(1-H^{t}(y))(1-G^{t}(y))}.$
First, we consider the case where the variables $X_{i}$ and censoring mechanism $Y_{i}$ are both heavy-tailed. In later sections this restriction is relaxed. More specifically, we assume now that $F$ and $G$ belong to the Fréchet max-domain of attraction, corresponding to a positive tail parameter. This implies that $F,$ $G$, and hence $H$, are regularly varying cdfs with respective tail indices $1/\gamma_{F},$ $1/\gamma_{G}$ and $1/\gamma_{H},$ where by (2.2) we obtain
$1/\gamma_{H}=1/\gamma_{F}+1/\gamma_{G}.$
In this case, by Theorem 1.2.1, [15],
$1-F^{t}(x)=\frac{1-F(tx)}{1-F(t)}\to x^{-1/\gamma_{F}}\quad\text{ as }\quad t\to\infty,$ (2.12)
and similarly for $G$ and $H$. Note that the latter convergence is uniform in $x>1$ by Potter’s bounds (see, e.g., Proposition B.1.9, [15]).
The sequence of integers $k_{n}$ is an intermediate sequence if
$k_{n}/n\to 0\quad\text{and}\quad k_{n}\to\infty\quad\text{as}\quad n\to\infty.$
Standard results of extreme value theory indicate that for an intermediate sequence $k=k_{n}$, the normalized upper order statistics $\{Z_{n-k+i,n}/Z_{n-k,n}\}_{i=1}^{k}$ given $Z_{n-k,n}=t$ are distributed as a vector of order statistics $\{V^{\ast}_{i,k}\}_{i=1}^{k}$ of a sample $\{V^{\ast}_{i}\}_{i=1}^{k}$ having cdf $H^{t}$ that converges to the Pareto distribution with index $1/\gamma_{H}$ as $t\to\infty$. However, to study the distributional convergence to the Pareto law with the correct target index $1/\gamma_{F}$ it is convenient to consider a product-limit construction, as follows (see also [18] for a moment construction).
Definition 1.
The extreme Kaplan–Meier estimator (EKM) is given by
$\amsmathbb{F}_{k,n}(x)=1-\prod_{i=1}^{k}\left[1-\frac{\delta_{[n-i+1:n]}}{i}\right]^{I(Z_{n-i+1,n}/Z_{n-k,n}\leq x)}.$ (2.13)
3 Extreme Kaplan–Meier Integrals
The asymptotic behaviour of extreme Kaplan–Meier integrals is studied in this section, defined as integrals with respect to the EKM estimator:
$S_{k,n}(\varphi)=\int\varphi\,\mathop{}\!\mathrm{d}\amsmathbb{F}_{k,n}.$ (3.1)
For technical reasons, such as applying the dominated convergence theorem, we assume that $\varphi$ is an almost everywhere (with respect to Lebesgue measure, written as a.e. for brevity) continuous function on $[1,\infty)$ and that $F,G,$ and therefore $H$ are eventually continuous.
Definition 2.
A continuous and eventually monotone function $\bar{\varphi}:[1,\infty)\to\amsmathbb{R}_{+}$ such that $|\varphi|\leq\bar{\varphi}$, is called an envelope of $\varphi.$
In the sequel, whenever we impose conditions on an envelope, we tacitly assume that the envelope exists. Otherwise, the condition is violated. The use of envelopes is technical in nature, to avoid requiring continuity and eventual monotonicity directly on $\varphi$. For a simple $\varphi$ function, the corresponding envelope is usually itself.
The following condition serves as a building block towards the general case, and is initially imposed, though later removed.
Temporary Condition 1.
The function $\varphi$ satisfies
$\varphi(x)=0\>\>\text{ for all }\>\>x>T\>\>\text{ and some }\>\>T<\infty.$ (3.2)
Important issues when studying EKM integrals arise from the fact that the “sample” is no longer independent. As a preliminary step, let us consider the behaviour of the EKM integral (3.1) conditioning on the denominator, that is given $Z_{n-k,n}=t.$ In this case, by analogy with Lemma 3.4.1, [15], the joint distribution of the vector of normalized order statistics $(Z_{n-k+1,n}/Z_{n-k,n},\ldots,Z_{n,n}/Z_{n-k,n})$ given $Z_{n-k,n}=t$ equals the joint distribution of order statistics $(V^{\ast}_{1,k},\ldots,V^{\ast}_{k,k})$ corresponding to i.i.d. random variables $\{V^{\ast}_{i}\}_{i=1}^{k}$ with cdf $H^{t}.$ By the Skorokhod representation theorem (e.g., p. 70 in [7]), on a rich enough probability space, it is possible to select $\{V^{\ast}_{i}\}_{i=1}^{k}$ in such a way that given $Z_{n-k,n}=t$ it holds $Z_{n-k+i,n}/Z_{n-k,n}=V_{i,k}^{\ast}$ almost surely, $i=1,\ldots,k.$ The following notation is useful in the sequel.
Definition 3.
For the random vectors
$\bar{Z}^{\ast}_{k}=(Z^{\ast}_{1,k},\dots,Z^{\ast}_{k,k})=(Z_{n-k+1,n}/Z_{n-k,n},\ldots,Z_{n,n}/Z_{n-k,n}),$
an array of functions $(f_{k,n})_{k,n\in\amsmathbb{N}}$, the baseline random variable $Z_{n-k,n}$, and a deterministic sequence $a_{1},a_{2},\ldots$ we say that $f_{k,n}(\bar{Z}^{\ast}_{k})$ is $o_{{P}}(a_{k})$ uniformly in large $t$ w.r.t. $Z_{n-k,n}$ if for every $\varepsilon>0$,
$\lim_{n/k,\,k\to\infty}\limsup\limits_{t\rightarrow\infty}{\amsmathbb{P}}(|f_{k,n}(\bar{Z}^{\ast}_{k})/a_{k}|>\varepsilon|\,Z_{n-k,n}=t)=0.$ (3.3)
We write in short (when no confusion arises) that $f_{k,n}(\bar{Z}^{\ast}_{k})=\overline{o}_{P}(a_{k})$. Analogously $f_{k,n}(\bar{Z}^{\ast}_{k})=\overline{O}_{P}(a_{k})$ using the notions of bounded in probability in place of convergence to zero in probability.
Note that $\bar{Z}^{\ast}_{k}\stackrel{{\scriptstyle d}}{{=}}\bar{V}^{\ast}_{k}$ given $Z_{n-k,n}=t.$ In the proofs, we work with $\bar{V}^{\ast}_{k}$ (considering its Skorokhod version, i.e. $\bar{V}^{\ast}_{k}=\bar{Z}^{\ast}_{k}$ a.s. given $Z_{n-k,n}=t$), the definitions of $\overline{o}_{P}$ and $\overline{O}_{P}$ can be easily recast in terms of $\bar{V}^{\ast}_{k}.$
We now provide the extreme value analogue to Theorem 1.1 of [38], where the decomposition is now in terms of sums of functions of upper order statistics, and the statement crucially involves the above-defined concept of uniform convergence in probability.
Theorem 3.1 (Decomposition, truncated case).
Let $\varphi$ have envelope $\bar{\varphi}$ and satisfy Temporary Condition 1. Let $k=k_{n}$ be an intermediate sequence. Then the following random sequence is $\overline{o}_{P}(k^{-1/2})$:
$r_{k,n}=S_{k,n}(\varphi)-\Big{\{}\frac{1}{k}\sum_{i=1}^{k}\varphi(Z_{i,k}^{\ast})\gamma_{0}^{t}(Z_{i,k}^{\ast})\delta^{\ast}_{[i:k]}+\frac{1}{k}\sum_{i=1}^{k}\gamma_{1}^{t}(Z_{i,k}^{\ast})(1-\delta^{\ast}_{[i:k]})-\frac{1}{k}\sum_{i=1}^{k}\gamma_{2}^{t}(Z_{i,k}^{\ast})\Big{\}},$ (3.4)
where $\{\delta^{\ast}_{i}\}_{i=1}^{k}$ are such that their concomitants $\{\delta^{\ast}_{[i:k]}\}_{i=1}^{k}$ equal $\{\delta_{[n-k+i:n]}\}_{i=1}^{k}$ a.s.
We proceed to state a general version of Theorem 3.1 by lifting the truncation condition (3.2). In the proof of this extension, we require analogues of (1.5) and (1.6) in [38], which help to control the tail. For that purpose, we impose the following key condition, which is natural for the limit variance to exist.
Condition 1.
For some $\varepsilon>0$
$\int_{1}^{\infty}|\bar{\varphi}(x)|^{2+\varepsilon}x^{\alpha(\varepsilon)}\mathop{}\!\mathrm{d}x<\infty,$ (3.5)
where $\alpha(\varepsilon)=1/\gamma_{G}-1/\gamma_{F}-1+\varepsilon.$
Proposition 3.2.
Assume Condition 1. Then for large $t$
$\int_{1}^{\infty}\varphi^{2}(\gamma^{t}_{0})^{2}\mathop{}\!\mathrm{d}H^{1,t}<\infty$ (3.6)
and
$\int_{1}^{\infty}|\varphi|\sqrt{C^{t}}\mathop{}\!\mathrm{d}F^{t}<\infty.$ (3.7)
Theorem 3.3 (Decomposition, general case).
Assume Condition 1, and let $k=k_{n}$ be an intermediate sequence. Then the random sequence $r_{k,n}$ defined by (3.4) is $\overline{o}_{P}(k^{-1/2})$.
The above representation, without additional assumptions allows us to prove that the extreme Kaplan–Meier integral $S_{k,n}(\varphi)$ (3.1) is a consistent estimator for
$S_{\circ}(\varphi):=\int_{1}^{\infty}\varphi\mathop{}\!\mathrm{d}F^{\circ}$
where $F^{\circ}(x)=(1-x^{-1/\gamma_{F}})I(x\geq 1).$
Theorem 3.4 (Weak consistency of EKM integrals).
Assume the conditions of Theorem 3.3. Then
$S_{k,n}(\varphi)\stackrel{{\scriptstyle P}}{{\to}}S_{\circ}(\varphi).$
Remark 3.5.
Weak consistency of extreme Kaplan–Meier integrals under weaker assumptions than Condition 1, may be possible with an approach as in Theorem 1.1 in [39], where the existence of the first moment is assumed for consistency (to the possibly modified target cdf) of the standard Kaplan–Meier integral. However, the proof of the latter result uses some specialized techniques which fall far from our setup, in the like of reverse-time supermartingales. In order not to further overload the notation and proofs, and to keep the narrative coherent, we simply continue to use the natural Condition 1 in Theorem 3.4.
The next theorem states the asymptotic normality of EKM integrals under the weak assumptions of Theorem 3.3. However, note that the centering sequence is random.
[…]
Define the random variable
$W^{\circ}(\varphi)=\varphi(V^{\circ})\gamma_{0}^{\circ}(V^{\circ})\delta^{\circ}+\gamma_{1}^{\circ}(V^{\circ})(1-\delta^{\circ})-\gamma_{2}^{\circ}(V^{\circ}).$ (3.8)
The variance of the random variable $W^{\circ}(\varphi)$ turns out to be the correct asymptotic variance in all of our weak convergence results.
Example 1.
Consider $\varphi(x)=\log(x)$. Then
$\gamma_{0}^{\circ}(V^{\circ})$ $=(V^{\circ})^{1/\gamma_{G}},\quad\gamma_{1}^{\circ}(V^{\circ})=(V^{\circ})^{1/\gamma_{G}}(\log(V^{\circ})+\gamma_{F}),$
$\gamma_{2}^{\circ}(V^{\circ})$ $=(V^{\circ})^{1/\gamma_{G}}(\log(V^{\circ})+\gamma_{F}-\gamma_{G})+\gamma_{G}-\gamma_{F},$
and so
$W^{\circ}(\varphi)=(V^{\circ})^{1/\gamma_{G}}(\gamma_{G}-\delta^{\circ}\gamma_{F})+\gamma_{F}-\gamma_{G}.$
Similarly, for $\varphi(x)=x$ we get
$W^{\circ}(\varphi)=(V^{\circ})^{1+1/\gamma_{G}}\Big{(}\delta^{\circ}+\frac{1-\delta^{\circ}}{1-\gamma_{F}}-\frac{1}{(1-\gamma_{F})(1+\gamma_{G})}\Big{)},$
and for $\varphi(x)=1\{x\leq x_{0}\}$ we obtain
$W^{\circ}(\varphi)$ $=1\{V^{\circ}\leq x_{0}\}\Big{(}1+\frac{\gamma_{H}}{\gamma_{G}}x_{0}^{-1/\gamma_{F}}((V^{\circ})^{1/\gamma_{H}}-1)-(V^{\circ})^{1/\gamma_{H}}x_{0}^{-1/\gamma_{F}}(1-\delta^{\circ})\Big{)}$
$\>\>+1\{V^{\circ}>x_{0}\}\Big{(}x_{0}^{1/\gamma_{G}}\frac{\gamma_{H}}{\gamma_{F}}+x^{-1/\gamma_{F}}\frac{\gamma_{H}}{\gamma_{G}}-1\Big{)}.$
Theorem 3.6.
Assume the conditions of Theorem 3.3. Then
$\sqrt{k}\int\varphi\,\mathop{}\!\mathrm{d}(\amsmathbb{F}_{k,n}-F^{Z_{n-k,n}})\stackrel{{\scriptstyle d}}{{\to}}N(0,\sigma^{2}_{\varphi}),$ (3.9)
where $\sigma^{2}_{\varphi}={\rm var}(W^{\circ}(\varphi)).$
In the next result, we replace in (3.9) the random centering sequence $\int\varphi\mathop{}\!\mathrm{d}F^{Z_{n-k,n}}$ with its nonrandom counterpart $\int\varphi\mathop{}\!\mathrm{d}F^{U_{H}(n/k)}$, where $U_{H}$ is the tail quantile function of $H,$ i.e.
$U_{H}(x)=\inf\{y:H(y)\geq 1-1/x\},\quad x>1.$
For this purpose, we need to impose the following condition on the tail quantile functions $U_{F}$ and $U_{H}:$ let a tail quantile function $U$ possess a derivative $U^{\prime}$ and let the following von Mises’ convergence hold
$\lim_{t\to\infty}\frac{tU^{\prime}(t)}{U(t)}=\gamma,$ (3.10)
which in itself is sufficient for $U$ to be regularly varying with tail index $\gamma$, and is standard in extreme value theory (see e.g. Theorem 1.1.11 and Corollary 1.1.12 in [15]). Notice that the condition (3.10) is equivalent to the following
$\lim_{t\to\infty}\frac{tK^{\prime}(t)}{1-K(t)}=\frac{1}{\gamma},$ (3.11)
where $U$ is the tail quantile function of cdf $K.$ Also, we need the following technical condition.
Condition 2.
Let $\varphi$ have a derivative $\varphi^{\prime}$ such that for some $\varepsilon^{\prime}>0$
$\int_{1}^{\infty}\bar{\varphi^{\prime}}(x)\,x^{-1/\gamma_{F}+\varepsilon^{\prime}}\mathop{}\!\mathrm{d}x<\infty.$ (3.12)
Hereinafter $\bar{\varphi^{\prime}}$ denotes an envelope of $\varphi^{\prime}.$ If the behaviour of $\varphi^{\prime}$ is regular enough, in particular, if $\varphi^{\prime}$ is continuous, then Condition 2 does not pose any additional complications, since it then follows from Condition 1.
Theorem 3.7.
Assume the conditions of Theorem 3.3. Furthermore, assume $U_{F}$ and $U_{H}$ satisfy (3.10), and $\varphi$ satisfies Condition 2. Then
$\sqrt{k}\int\varphi\,\mathop{}\!\mathrm{d}(F^{Z_{n-k,n}}-F^{U_{H}(n/k)})\stackrel{{\scriptstyle P}}{{\to}}0.$ (3.13)
As an immediate consequence of Theorems 3.6 and 3.7 we get the following result.
Corollary 3.8.
Under the assumptions of Theorem 3.7, we have
$\sqrt{k}\int\varphi\,\mathop{}\!\mathrm{d}(\amsmathbb{F}_{k,n}-F^{U_{H}(n/k)})\stackrel{{\scriptstyle d}}{{\to}}N(0,\sigma^{2}_{\varphi}).$ (3.14)
The convergence result of Corollary 3.8 is of independent interest, since it is not always possible to replace a nonrandom centering sequence with a constant in limit relations appearing in extreme value theory. For instance, it is not possible in Smirnov’s lemma and its corollary regarding limit behavior of intermediate order statistics, see e.g. Theorems 2.2.1 and 2.4.1 in [15].
Nevertheless, such a non-random replacement is possible in our case if the intermediate sequence $k=k(n)$ does not grow too fast with respect to $n.$ The precise speed is determined through the quality of the regular variation property of $U,$ which in turn is governed by a second order parameter denoted by $\rho,$ see [33]. Thus, to control this speed it is necessary to impose the following standard second order condition.
Condition 3.
The tail quantile function of $F$ satisfies
$\lim_{t\to\infty}\frac{U_{F}(tx)/U_{F}(t)-x^{\gamma_{F}}}{a(t)}=x^{\gamma_{F}}h_{\rho}(x),\quad x\geq 1,$ (3.15)
for some $\rho\leq 0,$ $a(t)$ a positive or negative function, regularly varying with index $\rho,$ and satisfying $a(t)\to 0$ as $t\to\infty,$ and
$h_{\rho}(x)=\left\{\begin{array}[]{cc}\log x,&\text{if }\rho=0,\\ (x^{\rho}-1)/\rho,&\text{ otherwise}.\end{array}\right.$ (3.16)
Properties and possible choices for the function $a(t)$ may be found in Section 2.3, [15].
Theorem 3.9.
Assume the conditions of Theorem 3.3. Furthermore, assume $\varphi$ satisfies Condition 2 and $U_{F}$ satisfies the second order Condition 3. Let $\sqrt{k}\,a(c_{n})\to\lambda$ with finite $\lambda$ and $c_{n}=(1-F(U_{H}(n/k)))^{-1},$ and if $\lambda\neq 0$ assume additionally that $\varphi^{\prime}$ is a.e. continuous. Then
$\sqrt{k}\int\varphi\,\mathop{}\!\mathrm{d}(F^{Z_{n-k,n}}-F^{\circ})\stackrel{{\scriptstyle P}}{{\to}}\lambda C(\gamma_{F},\rho),$ (3.17)
where
$C(\gamma_{F},\rho)=\int_{1}^{\infty}x\varphi^{\prime}(x)h_{\rho}(x^{1/\gamma_{F}})F^{\circ}(\mathop{}\!\mathrm{d}x).$
As an immediate consequence of Theorems 3.6 and 3.9, we get the following result.
Corollary 3.10.
Under the assumptions of Theorem 3.9, we have
$\sqrt{k}\int\varphi\,\mathop{}\!\mathrm{d}(\amsmathbb{F}_{k,n}-F^{\circ})\stackrel{{\scriptstyle d}}{{\to}}N(\lambda C(\gamma_{F},\rho),\sigma^{2}_{\varphi}).$ (3.18)
Remark 3.11.
Notice that for Theorem 3.9 to hold, we have required Condition 1. In particular, if $\varphi$ grows slower than a power function, the condition $\gamma_{F}<\gamma_{G}$ implies the required integrability. This non-heavy censoring regime for asymptotic normality of censored extremes has been observed before for related estimators (related to $\varphi=\log$), see for instance Remark 1 in [6].
Remark 3.12.
Analysing the joint behaviour of empirical versions of $H^{0,t}$ and $H^{1,t}$ is a somewhat direct route to prove asymptotic normality of $S_{k,n}(\varphi)$ using the functional delta method. However, there is an important caveat in that such assertion can only be proved for differentiable $\varphi$ and integrals up to an endpoint $T$ with $F(T)<1$, which is equivalent to $\varphi$ satisfying Condition 1. The above approach allows us to establish asymptotic results on the maximal interval, which in turn allows integration of infinite-support $\varphi$ functions.
4 Residual estimators revisited
4.1 Generalized residual estimators
As before, let $X_{1},\dots,X_{n}$ denote an i.i.d. sequence of random variables with cdf $F,$ and $X_{1,n}\leq\ldots\leq X_{n,n}$ denote the order statistics of $\{X_{i}\}_{i=1}^{n}.$ Denote by $U_{F}(t),$ the tail quantile function of $F,$ i.e. $U_{F}(t)=\inf\{x:F(x)\leq 1-1/t\}.$ For when $F$ belongs to the Fréchet max-domain of attraction (write $F\in\mathcal{D}(G_{\gamma})$ with $\gamma>0$), [34] studied the limit behavior of statistics of the form
$R_{k,n}=\frac{1}{k}\sum_{i=1}^{k}\vartheta(X_{n-i+1,n}/X_{n-k,n}),$
where $\vartheta:[1,\infty)\to\amsmathbb{R}$ is some suitable function, which is here assumed to be almost everywhere continuous.
Theorem 4.1 (Theorem 4.1, [34]).
Let $|\vartheta(x)|\leq Ax^{c},$ $x\geq 1,$ for some $A>0$ and $c<(2\gamma)^{-1}.$ Let $\xi,\xi_{1},\xi_{2},\ldots$ be independent Pareto$(1)$ random variables such that $X_{i}=U_{F}(\xi_{i}),$ $i\geq 1.$ Define $\mu(t)=\amsmathbb{E}[\vartheta(U_{F}(\xi t)/U_{F}(t))I(U_{F}(t)>0)].$ Let $k=k_{n}$ be an intermediate sequence. Then
$\sqrt{k}\big{(}R_{k,n}-\mu(\xi_{n-k,n})\big{)}\stackrel{{\scriptstyle d}}{{\to}}N(0,{\rm var}[\vartheta(\xi^{\gamma})]).$ (4.1)
In further theorems 4.2 and 4.5 ibid., it was proven that $\mu(\xi_{n-k,n})$ in (4.1) can be replaced with the nonrandom limit $\amsmathbb{E}[\vartheta(\xi^{\gamma})]$ under some additional conditions imposed on $F,$ $\vartheta$ and $k.$
It turns out that the direct extension of Theorem 4.1 to the Gumbel and Weibull max-domains of attraction leads to a degenerate limit in (4.1) in almost all cases. Recall the notation $\tau_{F}=\inf\{x:F(x)=1\},$ the right endpoint of $F.$ Specifically, we have the following result, whose proof is analogous to the proof of Theorem 4.1, and thus omitted.
Theorem 4.2.
Let $F\in\mathcal{D}(G_{\gamma})$ with $\gamma\leq 0$ and $\tau_{F}>0.$ Let $\vartheta:[1,\infty)\to\amsmathbb{R}$ satisfy $|\vartheta(x)|\leq Ax^{c},$ $x\geq 1,$ for some $A>0$ and $c>0.$ Assume $\mu(t),$ $k$ and $\xi,\xi_{1},\xi_{2},\ldots$ are as in Theorem 4.1. Then
$\sqrt{k}\big{(}R_{k,n}-\mu(\xi_{n-k,n})\big{)}\stackrel{{\scriptstyle P}}{{\to}}0.$ (4.2)
A reason of a degenerate limit in (4.2) is that for $\gamma\leq 0$ and $\tau_{F}>0$ the ratio $U_{F}(xt)/U_{F}(t)$ tends to $1$ as $t\to\infty$ (which follows from techniques in [15]), see for details Theorem 4.4 and Corollary 4.6.
It follows that to derive a non-degenerate limit, a faster than $\sqrt{k}$ normalizing sequence is required in (4.2). Consequently, a downside of using a residual estimators technique to develop extreme value estimators (in presence of random censorship) is that we may be faced with situations where positive and negative values of the extreme value index are allowed, but the rate of convergence of the estimator inconveniently varies. Another disadvantage is that $\vartheta(1)$ does not depend on $\gamma<0.$ This makes it impossible to construct extreme value index estimators based solely on $R_{k,n}.$
For this reason, let us consider an alternative and general approach. As before, let $\xi,\xi_{1},\xi_{2},\ldots$ be independent Pareto$(1)$ random variables with $n$-th order statistics denoted by $\xi_{1,n}\leq\ldots\leq\xi_{n,n}.$
Definition 4.
Let $f:[1,\infty)\times[1,\infty)\to[0,\infty)$, and $g,\,g_{1},\,g_{2}:[1,\infty)\to[0,\infty)$ be measurable functions such that for all $x$
$f(x,t)\to g(x),\quad t\to\infty,$ (4.3)
and for all $x$ there exists $T>0$, such that for $t\geq T$
$g_{1}(x)\leq f(x,t)\leq g_{2}(x).$ (4.4)
Let $\vartheta:[0,\infty)\to\amsmathbb{R}$. We are interested in the limit behavior of the following statistic
$R^{f}_{k,n}=\frac{1}{k}\sum_{j=1}^{k}\vartheta\left(f\Big{(}\frac{\xi_{n-j+1,n}}{\xi_{n-k,n}},\xi_{n-k,n}\Big{)}\right),$
where $k=k_{n}$ is an intermediate sequence.
Remark 4.3.
For technical reasons we generally assume that $g_{1},\,g_{2}$ can be modified to tend to $\infty$ together with $f(x,t)$. However, if $\bar{\vartheta}$ and $\bar{\vartheta^{\prime}}$ are eventually non-decreasing (resp. non-increasing), then it suffices to require $g_{2}$ (resp. $g_{1}$) to tend to $\infty$ together with $f(x,t).$
For the asymptotic results we require the following quantity:
$\mu_{f}(t)=\amsmathbb{E}\big{[}\vartheta(f(\xi,t))\big{]}.$ (4.5)
Theorem 4.4.
Let $k=k_{n}$ be an intermediate sequence, and assume for some $\delta>0$
$\amsmathbb{E}\Big{[}\max\big{\{}\bar{\vartheta}(g_{1}(\xi)),\bar{\vartheta}(g_{2}(\xi))\big{\}}\Big{]}^{1+\delta}<\infty.$ (4.6)
Then $R^{f}_{k,n}$ is a consistent estimator for $\amsmathbb{E}[\vartheta(g(\xi))].$ If, moreover, (4.6) is satisfied for some $\delta>1,$ then
$\sqrt{k}\big{(}R^{f}_{k,n}-\mu_{f}(\xi_{n-k,n})\big{)}\stackrel{{\scriptstyle d}}{{\to}}N\big{(}0,{\rm var}[\vartheta(g(\xi))]\big{)}.$ (4.7)
To replace $\mu_{f}(\xi_{n-k,n})$ with $\mu_{f}(\infty):=\amsmathbb{E}[\vartheta(g(\xi))]$ in Theorem 4.4, we need to impose another condition (compare with Condition 3).
Condition 4.
The following second-order type condition holds
$\frac{f(x,t)-g(x)}{\hat{a}(t)}\to\tilde{g}(x),\quad t\to\infty,$ (4.8)
where $\hat{a}$ is some positive or negative function. Furthermore, there exists a measurable function $\hat{g}$ such that for all $x$ and $t\geq T$
$|f(x,t)-g(x)|\leq|\hat{a}(t)|\hat{g}(x).$ (4.9)
In particular, if $\tilde{g}$ is regularly varying, then (4.8) implies (4.9).
Theorem 4.5.
Let the conditions of Theorem 4.4 and Condition 4 hold. Assume $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\lambda$ with finite $\lambda,$ $\vartheta$ has a derivative ϑ,superscriptitalic-$\vartheta^{\prime},$ and ϑsuperscriptitalic-$\vartheta^{\prime}$ is almost everywhere continuous if $\lambda\neq 0.$ Assume also
$\amsmathbb{E}\Big{[}\hat{g}(\xi)\max\big{\{}\bar{\vartheta^{\prime}}(g_{1}(\xi)),\bar{\vartheta^{\prime}}(g_{2}(\xi))\big{\}}\Big{]}<\infty.$ (4.10)
Then
$\sqrt{k}\big{(}R^{f}_{k,n}-\amsmathbb{E}[\vartheta(g(\xi))]\big{)}\stackrel{{\scriptstyle d}}{{\to}}N(C\,\lambda,{\rm var}[\vartheta(g(\xi))]\big{)},$ (4.11)
where $C=\amsmathbb{E}\big{[}\tilde{g}(\xi)\vartheta^{\prime}(g(\xi))\big{]}.$
4.2 Applications
In this section, we present some applications of Theorem 4.4 and Theorem 4.5. The most natural candidate function is
$f(x,t)=\frac{U_{F}(xt)}{U_{F}(t)}$
which appears in Theorem 4.1 and Theorem 4.2.
Thus, let $F\in\mathcal{D}(G_{\gamma})$ and assume for convenience that $\tau_{F}>0.$ By the remark below Theorem 4.2, $g(x)=x^{\gamma_{+}},$ where $\gamma_{+}=\max(\gamma,0)$. Moreover, by Potter’s bounds, for arbitrary $\delta_{1},\delta_{2}>0$, there exists $T=T(\delta_{1},\delta_{2})$ such that for $t\geq T$
$g_{1}(x):=\max\big{(}(1-\delta_{1})x^{\gamma_{+}-\delta_{2}}),1\big{)}\leq f(x,t)\leq(1+\delta_{1})x^{\gamma_{+}+\delta_{2}}=:g_{2}(x),$
where we select $\delta_{2}<\gamma$ for positive $\gamma.$
Next, if $\gamma>0$ then we can use the second order Condition 3. To derive an analogous condition in case $\gamma\leq 0,$ recall that $F\in\mathcal{D}(G_{\gamma})$ is equivalent to
$\lim_{t\to\infty}\frac{U_{F}(tx)-U_{F}(t)}{a(t)}=h_{\gamma}(x),\quad x>0,$ (4.12)
for some positive function $a,$ see, e.g., Theorem 1.1.6 in [15]. Thus, for $\gamma\leq 0$ and $x>0$
$\lim_{t\to\infty}\frac{U_{F}(tx)/U_{F}(t)-x^{\gamma_{+}}}{a(t)/U_{F}(t)}=h_{\gamma}(x),$ (4.13)
hence in (4.8) we can set
$\tilde{g}(x)=x^{\gamma_{+}}h_{\hat{\rho}}(x)$ (4.14)
with $\hat{\rho}=\rho I(\gamma>0)+\gamma I(\gamma\leq 0)$ and $\hat{a}$ equal to $a$ in Condition 3 for $\gamma>0$ and $a/U$ in (4.13) for $\gamma\leq 0,$ respectively. Next, the conditions of Lemma 4.6, [34], are satisfied under Condition 3 for $\gamma>0$ and (4.13) for $\gamma\leq 0,$ respectively, and thus in (4.9) we can set $\hat{g}(x)=Kx^{\gamma_{+}+\delta}$ for some $K>0,$ $\delta>0$ and $t\geq T.$ Finally, since $\hat{a}$ is regularly varying in our case, it suffices to require $\sqrt{k}\hat{a}(n/k)\to\lambda$ instead of $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\lambda$ by Smirnov’s lemma (see, e.g., Lemma 2.2.3 in [15] or (A.43)).
The following result is an immediate consequence of Theorem 4.4 and Theorem 4.5.
Corollary 4.6.
  1. (i)
    Let $F\in\mathcal{D}(G_{\gamma})$ with $\gamma\in\amsmathbb{R}$ and $\tau_{F}>0$, and $\vartheta:[1,\infty)\to\amsmathbb{R}$ satisfy $\amsmathbb{E}\big{[}(\bar{\vartheta}(\xi^{\gamma_{+}+\delta}))^{1+\delta}\big{]}<\infty$ for some $\delta>0,$
    and $\xi\sim\text{Pareto}(1).$ Assume $\mu(t),$ $k$ and $\xi_{1},\xi_{2},\ldots$ are as in Theorem 4.1. Then $R_{k,n}$ is consistent for $\amsmathbb{E}\big{[}\vartheta(\xi^{\gamma_{+}})\big{]}.$ If, moreover, $\amsmathbb{E}\big{[}(\bar{\vartheta}(\xi^{\gamma_{+}+\delta}))^{2+\delta}\big{]}<\infty$ for some $\delta>0,$ then
    $\sqrt{k}\big{(}R_{k,n}-\mu(\xi_{n-k,n})\big{)}\stackrel{{\scriptstyle d}}{{\to}}N(0,{\rm var}[\vartheta(\xi^{\gamma_{+}})]).$
  2. (ii)
    Assume additionally Condition 3 for $\gamma>0,$ $\sqrt{k}\hat{a}(n/k)\to\lambda$ with finite $\lambda$ for $\hat{a}$ defined above and almost everywhere continuity of $\vartheta^{\prime}$ if $\lambda\neq 0.$ Finally, assume $\amsmathbb{E}\big{[}\xi^{\gamma_{+}+\delta}\bar{\vartheta^{\prime}}(\xi^{\gamma_{+}+\delta})\big{]}<\infty$.
    Then
    $\sqrt{k}\big{(}R_{k,n}-\amsmathbb{E}\big{[}\vartheta(\xi^{\gamma_{+}})\big{]}\big{)}\stackrel{{\scriptstyle d}}{{\to}}N(C\lambda,{\rm var}[\vartheta(\xi^{\gamma_{+}})]),$
    where $C=\amsmathbb{E}\big{[}\tilde{g}(\xi)\vartheta^{\prime}(\xi^{\gamma_{+}})\big{]}$ and $\tilde{g}$ is defined by (4.14).
Remark 4.7.
In the above, it is always possible to find an intermediate sequence $k=k_{n}$ such that $\sqrt{k}\hat{a}(n/k)\to\lambda$ is satisfied. Indeed, if $\gamma>0$ then $\hat{a}(t)\to 0$ as $t\to\infty$ by definition (see Condition 3), and if $\gamma\leq 0$ then $\hat{a}(t)\to 0$ as $t\to\infty$ by Lemma 1.2.9, [15].
Note again that to estimate the extreme value index, Corollary 4.6 is not enough since for negative $\gamma$, $R_{k,n}$ is a consistent estimator of $\vartheta(1)$, which itself does not depend on $\gamma.$ Therefore, we now consider another function $f(x,t).$
As a starting point, note that it follows from (4.12) that
$\lim_{t\to\infty}\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)}=h_{\gamma_{-}}(x)$ (4.15)
with $\gamma_{-}=\min(\gamma,0)$ and the same $a(t)$ as in (4.12), see, e.g., formula (3.5.4) in [15]. The limit (4.15) implies that $\log U_{F}(t)$ is extended regularly varying with index $\gamma_{-},$ see the properties of such functions in section B.2, [15]. The relation (4.15) has previously been used to establish the asymptotic behavior of the moment estimator of EVI, see [17]. Thus, let us consider the candidate function
$f(x,t)=\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)},$
and therefore $g(x)=h_{\gamma_{-}}(x).$ For $g_{1}$ and $g_{2}$ in (4.4) one can use 00 and $x\mapsto\delta x^{\delta}$ for some $\delta\in(0,1)$ if $\gamma\geq 0$ by Corollary B.2.10, [15], and
$x\mapsto\gamma^{-1}\big{(}\max[(1\mp\delta)x^{\gamma\mp\delta}-(1\pm\delta),0]\big{)}$ (4.16)
for some $\delta\in(0,-\gamma),$ if $\gamma<0$ by Theorem B.2.2 ibid. and Potter’s bounds, respectively.
Next, the second order condition for $\log U_{F}(t)$ reads as follows
$\lim_{t\to\infty}\frac{\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)}-h_{\gamma_{-}}(x)}{Q(t)}=\int_{1}^{x}s^{\gamma_{-}-1}\int_{1}^{s}u^{\rho^{\prime}-1}\mathop{}\!\mathrm{d}u\mathop{}\!\mathrm{d}s=:H_{\gamma_{-},\rho^{\prime}}(x),$ (4.17)
for $x>0$, some $\rho^{\prime}\leq 0$ and $Q(t)$ positive or negative with $Q(t)\to 0,t\to\infty,$ see, e.g., formula (3.5.11) in [15]. Here, the function $H_{\gamma,\rho}$ can be re-expressed as
$H_{\gamma,\rho}(x)=\frac{1}{\rho}\left(\frac{x^{\gamma+\rho}-1}{\gamma+\rho}-\frac{x^{\gamma}-1}{\gamma}\right),$ (4.18)
which for the cases $\gamma=0$ and $\rho=0$ is understood to be equal to the limit of (4.18) as $\gamma\to 0$ or $\rho\to 0,$ respectively. Note that (4.17) follows from the standard extreme value theory second order condition
$\lim_{t\to\infty}\frac{\frac{U_{F}(tx)-U_{F}(t)}{a(t)}-h_{\gamma}(x)}{A(t)}=H_{\gamma,\rho}(x),$ (4.19)
if $\gamma\neq\rho,$ see, e.g., formula (3.5.10) in [15], where the relation between $Q$ and $A$ is stated by formula (3.5.12) ibid. and $\rho^{\prime}$ is the function of $\gamma$ and $\rho$ defined in Lemma B.3.16 ibid.
Hence, (4.17) implies $\tilde{g}=H_{\gamma_{-},\rho^{\prime}}.$ Since $\gamma_{-}$ and $\rho^{\prime}$ are non-positive, one can derive similarly to Lemma 4.6, [34], for some $K>0$ and $\delta>0$
$\left|\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)}-h_{\gamma_{-}}(x)\right|\leq K|Q(t)|\,x^{\gamma_{-}+\delta},\quad x\geq 1,t\geq T.$ (4.20)
Clearly, $\delta$ in (4.20) can be chosen such that $\gamma+\delta<0$ for negative $\gamma.$ Finally, $Q(t)$ appearing in (4.17) is regularly varying with index $\rho^{\prime}$ by Theorem B.3.1, [15], therefore as above we can require $\sqrt{k}Q(n/k)\to\lambda$ instead of $\sqrt{k}Q(\xi_{n-k,n})\to\lambda.$
Consider the statistic
$T_{k,n}=\frac{1}{k}\sum_{i=1}^{k}\vartheta\left(\frac{\log X_{n-i+1,n}-\log X_{n-k,n}}{a(U^{-1}(X_{n-k,n}))/X_{n-k,n}}\right)$
for some function $\vartheta$ and denote
$\mu_{f}(t)=\amsmathbb{E}\,\vartheta\left(\frac{\log U_{F}(t\xi)-\log U_{F}(t)}{a(t)/U_{F}(t)}\right),$
where $\xi\sim\text{Pareto}(1)$ as before.
The following result is a consequence of Theorem 4.4 and Theorem 4.5.
Corollary 4.8.
  1. (i)
    Let $F\in\mathcal{D}(G_{\gamma})$ with $\gamma\in\amsmathbb{R}$ and $\tau_{F}>0.$ Let $\vartheta:[0,\infty)\to\amsmathbb{R},$ and if $\gamma\geq 0,$ satisfy $|\vartheta(x)|\leq A\max(x^{c},1),$ $x\geq 0,$ for some $A>0$ and $c>0.$ Assume $k$ and $\xi_{1},\xi_{2},\ldots$ are as in Theorem 4.1. Then $T_{k,n}$ is consistent for $\amsmathbb{E}\big{[}\vartheta(h_{\gamma_{-}}(\xi))\big{]}$ and
    $\sqrt{k}\big{(}T_{k,n}-\mu_{f}(\xi_{n-k,n})\big{)}\stackrel{{\scriptstyle d}}{{\to}}N\big{(}0,{\rm var}\big{[}\vartheta(h_{\gamma_{-}}(\xi))\big{]}\big{)}.$
  2. (ii)
    Assume additionally (4.17), $\sqrt{k}Q(n/k)\to\lambda$ with finite $\lambda$ for $Q$ defined in (4.17), $|\vartheta^{\prime}(0)|<\infty,$ and almost everywhere continuity of $\vartheta^{\prime}$ if $\lambda\neq 0.$ Finally, if $\gamma\geq 0,$ assume $|\vartheta^{\prime}(x)|\leq A\max(x^{c},1),$ $x\geq 0,$ for some $A>0$ and $c>0.$ Then
    $\sqrt{k}\big{(}T_{k,n}-\amsmathbb{E}\big{[}\vartheta(h_{\gamma_{-}}(\xi))\big{]}\big{)}\stackrel{{\scriptstyle d}}{{\to}}N\big{(}C\lambda,{\rm var}\big{[}\vartheta(h_{\gamma_{-}}(\xi))\big{]}\big{)},$
    where $C=\amsmathbb{E}\big{[}H_{\gamma_{-},\rho^{\prime}}(\xi)\vartheta^{\prime}(h_{\gamma_{-}}(\xi))\big{]}$ and $H_{\gamma_{-},\rho^{\prime}}$ is defined by (4.17).
4.3 Extension of the results of Section 3
In this section, we extend the results of Section 3 using ideas from Section 4.1. Let us use the same notation as in Section 3. Crucially, unlike in Section 3, we do not assume that $F,$ $G$ and $H$ are regularly varying distribution functions.
Let us consider a sequence $\xi,\xi_{1},\xi_{2},\ldots$ of independent Pareto$(1)$ random variables such that a.s. $Z_{i}=U_{H}(\xi_{i}),$ $i\geq 1,$ and a function $f(x,t)$ satisfying Definition 4. Let $\vartheta:[0,\infty)\to\amsmathbb{R}$ be a measurable function as in Section 4.1. The following extreme Kaplan-Meier integral is of our main interest:
$S_{k,n}^{f}(\vartheta)=\int\vartheta(f(x,\xi_{n-k,n}))\amsmathbb{F}^{\xi}_{k,n}(\mathop{}\!\mathrm{d}x),$ (4.21)
where $\amsmathbb{F}^{\xi}_{k,n}$ is the extreme Kaplan-Meier estimator built by the sequence $\{\xi_{i}\},$ i.e.
$\amsmathbb{F}_{k,n}^{\xi}(x)=1-\prod_{i=1}^{k}\left[1-\frac{\delta_{[n-i+1:n]}}{i}\right]^{I(\xi_{n-i+1,n}/\xi_{n-k,n}\leq x)}.$
Thus $S^{f}_{k,n}(\vartheta)$ can be thought as an extreme Kaplan-Meier integral (3.1) for the function $\vartheta^{\xi_{n-k,n}}_{f}(x)=\vartheta(f(x,\xi_{n-k,n}))$ built through the sequence $\{\xi_{i}\}.$ The only, but substantial, difference is that the function $\vartheta^{\xi_{n-k,n}}_{f},$ the analogue of $\varphi$ in (3.1), is random.
Since we use $\{\xi_{i}\}$ as $\{Z_{i}\}$ in (4.21), the cumulative distribution functions $\tilde{F}$ and $\tilde{G}$ of the analogues of $\{X_{i}\}$ and $\{Y_{i}\},$ respectively, should be determined as well. The correct choice is clearly
$\tilde{F}(x)=F(U_{H}(x)),\quad\tilde{G}(x)=G(U_{H}(x)),\quad x\geq 1,$
and, evidently, $\tilde{H}(x)=H(U_{H}(x))=1-1/x,$ $x\geq 1.$ Further, let us define $\tilde{H}^{i,t},$ $i=0,1,$ and $\tilde{\gamma}^{t}_{j},$ $j=0,1,2,$ similarly to the Section 3 by replacing $F^{t},$ $G^{t},$ $H^{t},$ and $\varphi$ with $\tilde{F}^{t},$ $\tilde{G}^{t},$ $\tilde{H}^{t},$ and $\vartheta^{t}_{f},$ respectively. In the same fashion, define the random variable
$\tilde{W}^{\circ}(\vartheta)=\vartheta(g(\xi))\tilde{\gamma}_{0}^{\circ}(\xi)\delta^{\circ}+\tilde{\gamma}_{1}^{\circ}(\xi)(1-\delta^{\circ})-\tilde{\gamma}_{2}^{\circ}(\xi).$ (4.22)
Note that the random variables $\{\delta^{\ast}_{i}\}$ and $\delta^{\circ}$ remain unchanged from Section 3. We introduce the following condition.
Condition 5.
Assume $1-\tilde{F}$ is a regularly varying function with index $-\alpha_{F},$ with $\alpha_{F}\geq 0.$ Assume also $\tilde{F}$ is differentiable and $\tilde{F}^{\prime}$ is eventually monotone. If $\alpha_{F}=0$ then assume additionally that $\tilde{F}^{\prime}$ is regularly varying.
Remark 4.9.
If $\vartheta$ is monotone and $f(x,t)$ is monotone in the first argument, then the conditions on eventual monotonicity of $\tilde{F}^{\prime}$ and even existence of $\tilde{F}^{\prime}$ can be omitted, see e.g. the proof of Theorem 5.1.
Note that Condition 5 implies that $1-\tilde{G}$ is also a regularly varying function because $1-\tilde{G}=(1-\tilde{H})/(1-\tilde{F})$ and $1-\tilde{H}(x)=1/x,\,x\geq 1$ Denote $\alpha_{G},$ the index of regular variation of $1-\tilde{G}.$ Note that $\alpha_{F}$ and $\alpha_{G}$ belong to the interval $[0,1],$ moreover, $\alpha_{F}+\alpha_{G}=1$ by the definition of $F,$ $G$ and $H$ and the properties of regularly varying functions. Given $F$ and $G$ are eventually continuous, it is not difficult to show that $\tilde{F}$ and $\tilde{G}$ are eventually continuous as well.
Example 2.
In this example, we show that the first part of Condition 5 is rather easily satisfied. In particular, it holds true if $F$ and $G$ belong to the same max-domain of attraction.
  1. 1.
    First, assume $\gamma_{F},\gamma_{G}>0.$ Then $H$ is a regularly varying function with index $-1/\gamma_{H}=-1/\gamma_{F}-1/\gamma_{G},$ and thus $U_{H}(y)$ is a regularly varying function with index $\gamma_{H}.$ By the properties of regularly varying functions (see e.g. Proposition 1.5.7 (ii) in [8]), $1-\tilde{F}$ is regularly varying function with index $-\alpha_{F}=-\gamma_{H}/\gamma_{F},$ and $\alpha_{F}>0.$
  2. 2.
    Next, assume $\gamma_{F},\gamma_{G}<0.$ In this case, $F$ and $G$ have the finite right endpoints $\tau_{F}$ and $\tau_{G},$ and $\tau_{H}=\min(\tau_{F},\tau_{G}).$ So, $F$ and $G$ can be represented as follows
    $1-F(\tau_{F}-x)=x^{-1/\gamma_{F}}\ell_{F}(x),\quad 1-G(\tau_{G}-x)=x^{-1/\gamma_{G}}\ell_{G}(x),\quad x\geq 0,$
    with $\ell_{F},\ell_{G}$ slowly varying at 0, and thus for $x\geq 0$
    $1-H(\tau_{H}-x)=\left\{\begin{array}[]{cc}x^{-1/\gamma_{F}-1/\gamma_{G}}\ell_{H}(x),&\tau_{F}=\tau_{G};\\ x^{-1/\gamma_{F}}\ell_{H}(x),&\tau_{F}<\tau_{G};\\ x^{-1/\gamma_{G}}\ell_{H}(x),&\tau_{F}>\tau_{G};\end{array}\right.$
    with possibly different slowly varying functions $\ell_{H}$ in each line. It follows that $H$ belongs to the Weibull max-domain of attraction, i.e. $\gamma_{H}$ is negative. Note that if $\tau_{F}>\tau_{G}$ ($\tau_{F}<\tau_{G}$) then $\gamma_{F}$ ($\gamma_{G}$) can be thought of as equal to $\infty$- ∞ and we can keep the convenient formula $1/\gamma_{H}=1/\gamma_{F}+1/\gamma_{G}.$ Based on this relation, we get
    $U_{H}(y)=\tau_{H}-y^{\gamma_{H}}\ell^{\sharp}_{H}(y),$
    with $\ell^{\sharp}_{H}$ slowly varying at infinity. Assume first $\tau_{F}\leq\tau_{G},$ then $\tau_{F}$ and $\tau_{H}$ coincide, and we derive by the properties of regularly varying functions
    $1-\tilde{F}(y)=1-F(\tau_{F}-y^{\gamma_{H}}\ell^{\sharp}_{H}(y))=y^{-\gamma_{H}/\gamma_{F}}\ell_{\tilde{F}}(y)$
    with some $\ell_{\tilde{F}}(y)$ slowly varying at infinity. Note that $\alpha_{F}=\gamma_{H}/\gamma_{F}$ is positive in this case. Next assume that $\tau_{F}>\tau_{G},$ then $\tau_{H}=\tau_{F}.$ In this case $\tilde{F}$ is not a proper cdf because it cannot exceed $F(\tau_{G})$. Indeed, we get
    $1-\tilde{F}(y)=1-F(\tau_{G}-y^{\gamma_{H}}\ell^{\sharp}_{H}(y))\to 1-F(\tau_{G})>0,\quad y\to\infty,$
    so actually $1-\tilde{F}$ is slowly varying at infinity and thus $\alpha_{F}=0.$
  3. 3.
    Finally, consider the case $\gamma_{F}=\gamma_{G}=0.$ It may be assumed that $\tau_{F}=\tau_{G}$ because other cases are degenerate as above. Then by the von Mises representation (see e.g. Theorem 1.2.6 in [15]), we get for $x<\tau_{F}$
    $1-H(x)$ $=$ $(1-F(x))(1-G(x))=\exp\left(-\int_{x_{0}}^{x}\frac{\mathop{}\!\mathrm{d}s}{f(s)}+c_{1}(x)\right)\exp\left(-\int_{x_{0}}^{x}\frac{\mathop{}\!\mathrm{d}s}{g(s)}+c_{2}(x)\right)$ (4.23)
    $=$ $\exp\left(-\int_{x_{0}}^{x}\Big{(}\frac{1}{f(s)}+\frac{1}{g(s)}\Big{)}\mathop{}\!\mathrm{d}s+c_{1}(x)+c_{2}(x)\right)$
    with $\lim_{x\uparrow\tau_{F}}c_{i}(x)=c_{i}>0,$ $i=1,2,$ $f,g$ positive, and $\lim_{x\uparrow\tau_{F}}f^{\prime}(x)=\lim_{x\uparrow\tau_{F}}g^{\prime}(x)=0$. Since $h=fg/(f+g)$ is such that $\lim_{x\uparrow\tau_{F}}h^{\prime}(x)=0,$ $H$ belongs to the Gumbel max-domain of attraction, i.e. $\gamma_{H}=0,$ by Theorem 1.2.6, [15].
    The neat formula $\alpha_{F}=\gamma_{H}/\gamma_{F}$ does no longer hold in this case since $\gamma_{F}=\gamma_{H}=0.$ Nevertheless, $\alpha_{F}$ can take both positive (if, e.g., $F$ and $G$ are exponential cdfs) and zero (if, e.g., $F$ is exponential cdf and $G$ is Gaussian cdf) values.
The following is a simple yet important result which complements Example 2, and is used in the proofs of the results from the next section.
Proposition 4.10.
Assume $F$ and $G$ belong to two possibly different max-domains of attraction, $1-\tilde{F}$ is a regularly varying function with index $-\alpha_{F},$ and $\alpha_{F}>0.$ Then $F$ and $H$ belong to the same max-domain of attraction and, moreover, $\tau_{F}=\tau_{H}.$
Finally, we introduce a second moment condition similar to Condition 1 and (4.6). Recall that $\bar{\vartheta}$ denotes an envelope of $\vartheta,$ and $g_{1},g_{2}$ are defined by Definition 4.
Condition 6.
There exists $\varepsilon>0$ such that
$\int_{1}^{\infty}\big{(}\max(\bar{\vartheta}(g_{1}(x)),\bar{\vartheta}(g_{2}(x)))\big{)}^{2+\varepsilon}x^{\tilde{\alpha}(\varepsilon)}\mathop{}\!\mathrm{d}x<\infty,$
where $\tilde{\alpha}(\varepsilon)=\alpha_{G}-\alpha_{F}-1+\varepsilon.$
Denote
$S_{\circ}^{f}(\vartheta):=\int_{1}^{\infty}\vartheta(g(y))\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}y)$
where similarly to Section 3, $\tilde{F}^{\circ}(x)=(1-x^{-\alpha_{F}})I(x\geq 1).$ If $\alpha_{F}=0$ then $\tilde{F}^{\circ}(x)$ should be understood as 00 for all $x\geq 1,$ and thus $S_{\circ}^{f}(\vartheta)=0$ as well. Clearly, the latter case is not useful.
Theorem 4.11.
Assume Condition 5 and Condition 6, let $k=k_{n}$ be an intermediate sequence. Then $S^{f}_{k,n}(\vartheta)$ is a consistent estimator for $S_{\circ}^{f}(\vartheta)$ and
$\sqrt{k}\big{(}S^{f}_{k,n}(\vartheta)-\mu_{f}(\xi_{n-k,n})\big{)}\stackrel{{\scriptstyle d}}{{\to}}N\big{(}0,{\rm var}[\tilde{W}^{\circ}(\vartheta)]\big{)},$ (4.24)
where
$\mu_{f}(t)=\int_{1}^{\infty}\vartheta(f(y,t))\tilde{F}^{t}(\mathop{}\!\mathrm{d}y).$
To replace $\mu_{f}(\xi_{n-k,n})$ with $\mu_{f}(\infty):=S_{\circ}^{f}(\vartheta)$ in the latter theorem, we impose a second-order type condition as was done in Theorem 3.9 and Theorem 4.5. However, in contrast to those theorems, the second-order behavior of both $f$ and $\tilde{F}^{t}$ impacts the asymptotics, and consequently we need to impose two second order conditions. First, keeping the notation of Section 4.1, we impose Condition 4 on $f$. Similar to Theorem 4.5, we need to control the behavior of $\vartheta^{\prime}$ motivating the following.
Condition 7.
Let $\vartheta$ have a derivative $\vartheta^{\prime}$ such that for some $\varepsilon^{\prime}>0$
$\int_{1}^{\infty}\hat{g}(x)\max\big{(}\bar{\vartheta^{\prime}}(g_{1}(x)),\bar{\vartheta^{\prime}}(g_{2}(x))\big{)}\,x^{-\alpha_{F}-1+\varepsilon^{\prime}}\mathop{}\!\mathrm{d}x<\infty.$
Secondly, similarly to Theorem 3.9, we require a second order condition on $U_{\tilde{F}},$ the analogue of Condition 3. Further we focus on the non-trivial case $\alpha_{F}>0$. The following analogue of Condition 2, handles the behavior of the derivative of the function $\vartheta_{g}=\vartheta\circ g.$
Condition 8.
Let $\vartheta_{g}$ have a derivative $\vartheta_{g}^{\prime}$ such that for some $\varepsilon^{\prime}>0$
$\int_{1}^{\infty}\bar{\vartheta_{g}^{\prime}}(x)\,x^{-\alpha_{F}+\varepsilon^{\prime}}\mathop{}\!\mathrm{d}x<\infty.$
Theorem 4.12.
Let the conditions of Theorem 4.11, Condition 4, Condition 7, and Condition 8 hold. Assume Condition 3 holds for $U_{\tilde{F}},$ $a=a_{\tilde{F}}$ and $\alpha_{F}>0$, that $\sqrt{k}a(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\lambda,$ $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\hat{\lambda}$ with finite $\lambda,$ $\hat{\lambda},$ and ϑsuperscriptitalic-$\vartheta^{\prime}$ and $\vartheta_{g}^{\prime}$ are almost everywhere continuous if $\hat{\lambda}\neq 0$ and $\lambda\neq 0,$ respectively. Then
$\sqrt{k}\big{(}S^{f}_{k,n}(\vartheta)-S_{\circ}^{f}(\vartheta)\big{)}\stackrel{{\scriptstyle d}}{{\to}}N(\mu,{\rm var}[\tilde{W}^{\circ}(\vartheta)]\big{)},$ (4.25)
where
$\mu=\hat{\lambda}\int\tilde{g}(x)\vartheta^{\prime}(g(x))\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}x)+\lambda\int x\vartheta_{g}^{\prime}(x)h_{\rho}(x^{\alpha_{F}})\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}x).$
Remark 4.13.
Note that in most practical cases only one term in asymptotic bias may be positive, since the auxiliary functions $a$ and $\hat{a}$ generally have different asymptotic behaviour. This is in line with classical results in statistics of extremes: depending on the extreme value index and the second order parameter, the formula for the asymptotic bias changes substantially, see e.g. Theorem 3.5.4, [15], and in particular formula (3.5.16) therein.
5 Extreme value index estimation
There are numerous extreme value index estimators proposed in the literature, see e.g. Section 6 in the review [23]. Residual estimators (see Section 4 for a formal definition) are used in some of them. For instance, the moment estimator, [17], is defined as follows
$g_{k,n}=M_{k,n}^{(1)}+1-\frac{1}{2}\left(1-\frac{\big{(}M_{k,n}^{(1)}\big{)}^{2}}{M_{k,n}^{(2)}}\right)^{-1},$ (5.1)
where
$M_{k,n}^{(\ell)}=\frac{1}{k}\sum_{i=1}^{k}\big{(}\log X_{n-i+1,n}-\log X_{n-k,n}\big{)}^{\ell}=\frac{1}{k}\sum_{i=1}^{k}f_{\ell}(X_{n-i+1,n}/X_{n-k,n}),\>\>\>\ell=1,2,$
are the residual estimators with $f_{\ell}(x)=(\log x)^{\ell}.$ The estimator $M_{k,n}^{(1)}$ was proposed in [28] for positive tail indices, while $g_{k,n}$ targets all max-domains of attraction. Note that $M_{k,n}^{(\ell)}$ can be rewritten as follows:
$M_{k,n}^{(\ell)}=\int_{1}^{\infty}f_{\ell}(x)F_{k,n}(\mathop{}\!\mathrm{d}x),$ (5.2)
where
$F_{k,n}(x)=\frac{1}{k}\sum_{i=1}^{k}I(X_{n-i+1,n}/X_{n-k,n}\leq x),$
is the empirical distribution function of the set of the normalized largest order statistics
$\{X_{n-i+1,n}/X_{n-k,n}\}_{i=1}^{k}.$
The aim of this section is to apply the ideas of previous sections for extreme value index estimation in the presence of random censorship. Specifically, we propose a version of the moment estimator (5.1), where we use the extreme Kaplan–Meier estimator $\amsmathbb{F}_{k,n}$ in place of $F_{k,n}$. Thus, denote
$\amsmathbb{M}_{k,n}^{(\ell)}=\int_{1}^{\infty}(\log x)^{\ell}\,\amsmathbb{F}_{k,n}(\mathop{}\!\mathrm{d}x),\>\>\>\ell=1,2,$ (5.3)
and then define $\mathbb{g}_{k,n}$ by replacing $M_{k,n}^{\ell}$ with $\amsmathbb{M}_{k,n}^{\ell},$ $\ell=1,2,$ in (5.1), that is
$\mathbb{g}_{k,n}=\amsmathbb{M}_{k,n}^{(1)}+1-\frac{1}{2}\left(1-\frac{\big{(}\amsmathbb{M}_{k,n}^{(1)}\big{)}^{2}}{\amsmathbb{M}_{k,n}^{(2)}}\right)^{-1}.$ (5.4)
In what follows we show the asymptotic properties of the resulting estimator $\mathbb{g}_{k,n}$ taking full advantage of the theory of extreme Kaplan–Meier integrals.
As in Condition 5, we assume that $1-F(U_{H}(x))$ and therefore $1-G(U_{H}(x))$ are regularly varying functions with indices $-\alpha_{F}$ and $-\alpha_{G},$ respectively. Recall that by definition of $F,$ $G$ and $H,$ $\alpha_{F},\alpha_{G}\in[0,1]$. Example 2 implies that this condition includes Case 1 and Case 3 considered by [18]:
  1. Case 1.
    $\gamma_{F}>0,$ $\gamma_{G}>0;$
  2. Case 2.
    $\gamma_{F}=\gamma_{G}=0$ and $x_{F}^{\ast}=x_{G}^{\ast}=\infty;$
  3. Case 3.
    $\gamma_{F}<0,$ $\gamma_{G}<0$ and $x_{F}^{\ast}=x_{G}^{\ast}.$
In [18] it is stated that other possibilities are not very interesting since they are close to either the “uncensored case” (when $\gamma_{F}<0$ and $\gamma_{G}>0$) or to the “completely censored situation” (when $\gamma_{F}>0$ and $\gamma_{G}<0$). However, the situation $\gamma_{F}=\gamma_{G}=0$ and $x_{F}^{\ast}=x_{G}^{\ast}<\infty$ does not fit the description and is not treated. Consequently, we continue with our own general condition.
The following result is the direct consequence of Theorem 3.4 and the first statement of Theorem 4.11.
Theorem 5.1.
Assume $F$ and $G$ belong to (possibly different) max-domains of attraction such that $1-F(U_{H}(x))$ is regularly varying function with index $-\alpha_{F},$ where $\alpha_{F}>1/2$. Additionally, assume $x^{\ast}_{F}>0.$ Let $k=k_{n}$ be an intermediate sequence. Then
$\mathbb{g}_{k,n}\stackrel{{\scriptstyle P}}{{\to}}\gamma_{F}.$
Notice that the positiveness of $\alpha_{F}$ in the latter theorem is crucial since in case of zero $\alpha_{F}$ the limit in probability of $\mathbb{g}_{k,n}$ does not depend on $\gamma_{F}.$ Moreover, this assumption ensures that $F$ and $H$ belong to the same maximum domain of attraction by Proposition 4.10.
If both $F$ and $G$ belong to the Fréchet or Weibull max-domains of attraction, we get by Example 2
$\alpha_{F}=\frac{\gamma_{H}}{\gamma_{F}}=\frac{\gamma_{G}}{\gamma_{F}+\gamma_{G}},$
hence for positive $\gamma_{F}$ and $\gamma_{G}$ the condition $\alpha_{F}>1/2$ is equivalent to the natural assumption $1/\gamma_{F}>1/\gamma_{G}$ (and $1/\gamma_{F}<1/\gamma_{G}$ for negative $\gamma_{F}$ and $\gamma_{G}$) appearing in [6], that is the tail of $F$ is lighter than the tail of $G.$
Now, let us consider the asymptotic normality of the estimator $\mathbb{g}_{k,n}.$ The proof of this property is based on Theorem 3.9 (for $\gamma_{F}>0$) and Theorem 4.11 (for $\gamma_{F}\leq 0$). We can derive this result immediately applying Theorem 4.12, see Corollary 5.4 below. Note that Theorem 4.12 requires imposing the second order conditions both on $U_{H}$ and $U_{\tilde{F}}.$ Our proof is slightly longer since we also show that it is possible to avoid such restriction for $\gamma_{F}\neq 0.$ For positive $\gamma_{F},$ we prove the asymptotic normality of the estimator $\mathbb{g}_{k,n}$ using the results of Section 3. For negative $\gamma_{F},$ using the regular variation property of $x_{F}^{\ast}-U_{F},$ we assume the analogue of the second order condition (4.19) with $a(t)=-\gamma_{F}(x^{\ast}_{F}-U_{F}(t)).$ Based on Remark B.3.15 in [15], let us consider the following condition
$\lim_{t\to\infty}\frac{\frac{x_{F}^{\ast}-U_{F}(tx)}{x^{\ast}_{F}-U_{F}(t)}-x^{\gamma_{F}}}{A(t)}=x^{\gamma_{F}}h_{\rho}(x)$ (5.5)
for some $\rho\leq 0,$ and $A(t)$ positive or negative, regularly varying with index $\rho$ and satisfying $A(t)\to 0$ as $t\to\infty,$ which is the analogue of Condition 3 for negative $\gamma_{F}.$
Recall the notation $c_{n}=(1-F(U_{H}(n/k)))^{-1}.$
Theorem 5.2.
Assume the conditions of Theorem 5.1.
  1. (i)
    Let $\gamma_{F}>0,$ $U_{F}$ satisfy the second order Condition 3 and $\sqrt{k}\,a(c_{n})\to\lambda$ with finite $\lambda.$ Then
[…]
    where
    $b=\lambda\frac{\gamma_{F}-\gamma_{F}\rho+\rho}{\gamma_{F}(1-\rho)^{2}},\quad\sigma^{2}=\frac{\gamma^{3}_{G}+\gamma_{G}\gamma_{F}^{2}(\gamma_{G}-\gamma_{F}+1)^{2}}{(\gamma_{G}-\gamma_{F})^{3}}.$ (5.7)
  2. (ii)
    Let $\gamma_{F}=0,$ $U_{H}$ and $U_{\tilde{F}}$ satisfy (4.17) and Condition 3 with auxiliary functions $Q,$ $\tilde{a}$ and second order parameters $\rho^{\prime},$ $\tilde{\rho},$ respectively. Assume $\sqrt{k}Q(n/k)\to\lambda$ and $\sqrt{k}\tilde{a}(n/k)\to\tilde{\lambda}$ with finite $\lambda,\tilde{\lambda}.$ Then (5.6) holds with
    $b=\lambda\frac{1+\alpha_{F}}{\alpha_{F}^{2}}+\tilde{\lambda}\frac{1-\tilde{\rho}+\alpha_{F}\tilde{\rho}}{(1-\tilde{\rho})^{2}},\quad\sigma^{2}=\frac{\alpha^{4}_{F}+(\alpha_{F}-\alpha_{G}+\alpha_{F}\alpha_{G})^{2}I(\alpha_{G}>0)}{\alpha_{F}(\alpha_{F}-\alpha_{G})^{3}}.$
  3. (iii)
    Let $\gamma_{F}<0,$ $U_{F}$ satisfy the second order condition (5.5), $\sqrt{k}A(c_{n})\to\lambda$ and $\sqrt{k}(1-U_{H}(n/k)/\tau_{H})\to\hat{\lambda}$ with finite $\lambda,\hat{\lambda}.$ Then (5.6) holds with
    $b$ $=$ $\lambda\frac{(1-\gamma_{F})(1-2\gamma_{F})}{1-\gamma_{F}-\rho}\left(\frac{2-3\gamma_{F}-\rho}{1-2\gamma_{F}-\rho}-\frac{2}{\gamma_{F}}+\frac{1}{\gamma_{H}(1-\gamma_{F})^{2}(1-2\gamma_{F})}\right)$
    $+\hat{\lambda}\left(\frac{(1-\gamma_{F})(1-\gamma_{F}-\gamma_{F}^{2})}{1-3\gamma_{F}}-\frac{\gamma_{F}}{2\gamma_{H}(1-\gamma_{F})(1-2\gamma_{F})}\right),$
    and $\sigma^{2}$ defined by (C.9).
Note that it is possible to prove the assertion $(i)$ of the latter theorem by imposing the second condition (4.19) instead of Condition 3, as it was done, e.g., in Theorem 3.5.4, [15], proving asymptotic normality of the moment estimator in the uncensored case. We avoid taking such a route to highlight that in case $\gamma_{F}>0$ the proof quite simply follows from Theorem 3.9. The slight difference in asymptotic biases between Theorem 5.2 $(i)$ and the corresponding case in Theorem 3.5.4, [15] is caused by different second order conditions and in particular different auxiliary functions. Moreover, setting $\gamma_{G}=\infty,$ $\gamma_{G}=-\infty$ and $\alpha_{G}=0$ for positive, negative and zero $\gamma_{F},$ respectively (corresponding to the non-censored case) the asymptotic variances in Theorem 5.2 coincide with all the corresponding asymptotic variances in Theorem 3.5.4, [15].
Remark 5.3 (Comparison with the moment estimator adapted for censoring).
The moment estimator adapted for censoring presented in [18] is based on the simple observation that $p=\gamma_{H}/\gamma_{F}$ and thus one can estimate $\gamma_{H}$ and $p$ to estimate $\gamma_{F}$ indirectly, and their conditions are of very different nature to the ones we impose. The asymptotic variance of their estimator is usually smaller than the asymptotic variance of $\mathbb{g}_{k,n}$. However, the method proposed in [18] works exclusively for the estimation of extreme value index. Moreover, the conditions imposed in [18] on the (uniformity and speed of) convergence of $p(z)={\amsmathbb{P}}(\delta=1|Z=z)$ to $p$ (whose definition coincides with our $\alpha_{F}$ under differentiability of $F$ and $G$ – a condition we do not require) are very strong and virtually unverifiable. We believe our conditions are either standard or offer clear and verifiable alternatives.
The condition $\gamma_{G}>\gamma_{F}$ imposed in Theorem 5.1 and Theorem 5.2 can be overcome by constructing a moment estimator for the censoring mechanism and for the entire sample, and then using $\gamma_{F}=(\frac{1}{\gamma_{H}}-\frac{1}{\gamma_{G}})^{-1}$ to provide a finite asymptotic variance estimator of $\gamma_{F}$.
Finally, it is also of independent interest to obtain a version of Theorem 5.2 following directly from Theorem 4.12. As mentioned above, the asymptotic bias in this version differs from the asymptotic bias in Theorem 5.2 since the auxiliary functions appearing in the second-order conditions are different.
Corollary 5.4.
Assume the conditions of Theorem 5.1. Let $U_{H}$ and $U_{\tilde{F}}$ satisfy (4.17) and Condition 3 with auxiliary functions $Q,$ $\tilde{a}$ and second order parameters $\rho^{\prime},$ $\tilde{\rho},$ respectively. Assume $\sqrt{k}Q(n/k)\to\lambda$ and $\sqrt{k}\tilde{a}(n/k)\to\tilde{\lambda}$ with finite $\lambda,\tilde{\lambda}.$ Suppose also $\gamma_{F}\neq\rho$ with $\rho$ defined by (4.19). Then (5.6) holds with the same $\sigma^{2}$ as in Theorem 5.2 and
$b$ $=$ $\lambda\frac{\alpha_{F}(1-\gamma_{F}^{-})^{2}(1-2\gamma_{F}^{-})+(1-2\gamma_{F}^{-}-\hat{\rho})}{\alpha_{F}^{2}(1-\gamma_{F}^{-})(1-\gamma_{F}^{-}-\hat{\rho})(1-2\gamma_{F}^{-}-\hat{\rho})}$
$+\tilde{\lambda}\frac{\alpha_{F}(1-\gamma_{F}^{-})(1-2\gamma_{F}^{-})(3\gamma_{F}^{-}+\tilde{\rho}-4(\gamma_{F}^{-})^{2}-2\gamma_{F}^{-}\tilde{\rho})+(1-2\gamma_{F}^{-}-\tilde{\rho})}{(1-\gamma_{F}^{-}-\tilde{\rho})(1-2\gamma_{F}^{-}-\tilde{\rho})},$
where $\hat{\rho}=\rho^{\prime}/\alpha_{F}$ and $\gamma_{F}^{-}=\min(\gamma_{F},0).$
Note that for $\gamma_{F}^{-}=0$ the bias in the latter theorem coincides with the bias in Theorem 5.2 (ii), since then $\rho^{\prime}=0$ by Lemma B.3.16 in [15].
Appendix A Proofs of Section 3
Proof of Theorem 3.1.
As shown above, we can find i.i.d. random variables $\{V^{\ast}_{i}\}_{i=1}^{k}$ such that the joint distribution of the set of their order statistics $(V^{\ast}_{1,k},\ldots,V^{\ast}_{k,k})$ equals the joint distribution of the set $(Z_{n-k+1,n}/Z_{n-k,n},\ldots,Z_{n,n}/Z_{n-k,n})$ given $Z_{n-k,n}=t.$ But it is not evident whether the random variables $\{\delta^{\ast}_{i}\}_{i=1}^{k}$ introduced in Theorem 3.1 are i.i.d. as well. The following lemma answers this question.
Lemma A.1.
The sequence of vectors $\{(V_{i}^{\ast},\delta^{\ast}_{i})\}_{i=1}^{k}$ is i.i.d.
Proof.
Since the distribution of $\{(Z_{i},\delta_{i})\}_{i=1}^{n}$ is completely defined by the distribution of $\{(X_{i},Y_{i})\}_{i=1}^{n},$ let us consider the latter. Let $\{T_{i}\}_{i=1}^{k}$ be the times of exceedances over $t$ by the sequence $\{Z_{i}\}_{i=1}^{n}$ given $Z_{n-k,n}=t,$ that is
$T_{i}=\min\{j>T_{i-1}:\;Z_{j}>t\},\quad T_{0}=0.$
Consider the joint distribution of the pairs $\{(X_{T_{i}},Y_{T_{i}})\}_{i=1}^{k}$ given $Z_{n-k,n}=t.$ First, the condition $Z_{n-k,n}=t$ means that $X_{T_{i}}\geq t,$ $Y_{T_{i}}\geq t,$ $i=1,\ldots,k,$ so it is easy to show that
$\mathcal{L}\left(\left.\binom{X_{T_{1}}}{Y_{T_{1}}},\ldots,\binom{X_{T_{k}}}{Y_{T_{k}}}\right|Z_{n-k,n}=t\right)=$
$\mathcal{L}\left(\left.\binom{X_{T_{1}}}{Y_{T_{1}}},\ldots,\binom{X_{T_{k}}}{Y_{T_{k}}}\right|\binom{\text{ the number of exceedances is exactly }k}{\text{and }X_{T_{i}}\geq t,Y_{T_{i}}\geq t,i=1,\ldots,k}\right)$
Clearly, the vector $(T_{1},\ldots,T_{k})$ is uniformly distributed on the set of all k-tuples of $\{1,\ldots,n\},$ thus the latter distribution is equal to
$\mathcal{L}\left(\left.\binom{X_{1}}{Y_{1}},\ldots,\binom{X_{k}}{Y_{k}}\right|\binom{\text{ the number of exceedances is exactly }k}{\text{and }X_{i}\geq t,Y_{i}\geq t,i=1,\ldots,k}\right).$
The latter distribution is clearly independent of $n,$ thus it is equal to
$\mathcal{L}\left(\left.\binom{X_{1}}{Y_{1}},\ldots,\binom{X_{k}}{Y_{k}}\right|X_{i}\geq t,Y_{i}\geq t,i=1,\ldots,k\right).$
Since $\{X_{i}\}$ and $\{Y_{i}\}$ are independent, this distribution coincides with
$\mathcal{L}\left(\binom{X^{\ast}_{1}}{Y^{\ast}_{1}},\ldots,\binom{X^{\ast}_{k}}{Y^{\ast}_{k}}\right),$
where $\{X_{i}^{\ast}\}_{i=1}^{k}$ and $\{Y_{i}^{\ast}\}_{i=1}^{k}$ are two independent i.i.d. samples with cdfs $F^{t}(x/t)$ and $G^{t}(x/t),$ respectively. Since the distribution of $(Z_{i},\delta_{i})$ is completely determined by the distribution of $(X_{i},Y_{i}),$ we can select $V_{i}^{\ast}=\min(X_{i}^{\ast}/t,Y_{i}^{\ast}/t)$ and $\delta_{i}^{\ast}=I(X_{i}^{\ast}\leq Y_{i}^{\ast}).$
Consider now the extreme Kaplan–Meier integral $S_{k,n}(\varphi),$ defined by (3.1), given $Z_{n-k,n}=t.$ By Lemma A.1 and the Skorokhod representation, under this condition the extreme Kaplan–Meier estimator (2.13) takes the following form
$\amsmathbb{F}_{k,n}^{t}(x)=1-\prod_{i=1}^{k}\left[1-\frac{\delta^{\ast}_{[i:k]}}{k-i+1}\right]^{I(V^{\ast}_{i,k}\leq x)},$
where the latter relation holds almost surely, and replacing in the latter $k$ by $n,$ $\delta^{\ast}_{[i:k]}$ by $\delta_{[i:n]},$ and $V_{i,k}^{\ast}$ by $Z_{i,n}$ gives us exactly the classical Kaplan–Meier estimator (2.3). In particular, this implies that, for fixed $k$ and $n$, we may follow along the lines of the initial steps of the proof of Theorem 1.1 in [38] (specifically, compare the below formula with (2.1) therein) to obtain
$S_{k,n}(\varphi)$
$=\frac{1}{k}\sum_{i=1}^{k}\varphi(V_{i}^{\ast})\gamma^{t}_{0}(V_{i}^{\ast})\delta^{\ast}_{i}[1+B_{ik}^{t}+C_{ik}^{t}]+\frac{1}{k}\sum_{i=1}^{k}\frac{1}{2}\varphi(V_{i}^{\ast})\delta^{\ast}_{i}e^{\Delta_{i}^{t}}[B_{ik}^{t}+C_{ik}^{t}]^{2},$ (A.1)
where
$B_{ik}^{t}=k\int_{1}^{V_{i}^{\ast}-}\log\left[1+\frac{1}{k(1-H_{k}^{t}(z))}\right]H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)-\int_{1}^{V_{i}^{\ast}-}\frac{1}{1-H_{k}^{t}(z)}H^{0,t}_{k}(\mathop{}\!\mathrm{d}z),$ (A.2)
$C_{ik}^{t}=\int_{1}^{V_{i}^{\ast}-}\frac{1}{1-H_{k}^{t}(z)}H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)-\int_{1}^{V_{i}^{\ast}-}\frac{1}{1-H^{t}(z)}H^{0,t}(\mathop{}\!\mathrm{d}z),$ (A.3)
$\Delta_{i}^{t}$ is between the first term in the rhs of (A.2) and second term in the rhs of (A.3), and
$H^{t}_{k}(z)=\frac{1}{k}\sum_{i=1}^{k}I(V_{i}^{\ast}\leq z),\quad H^{j,t}_{k}(z)=\frac{1}{k}\sum_{i=1}^{k}I(V_{i}^{\ast}\leq z,\delta_{i}^{\ast}=j),\quad j=0,1.$
Moreover,
$\frac{1}{k}\sum_{i=1}^{k}\varphi(V_{i}^{\ast})\gamma^{t}_{0}(V_{i}^{\ast})\delta^{\ast}_{i}C_{ik}^{t}=$ $-\int\int\int\frac{I(v<u,v<w)\varphi(w)\gamma_{0}^{t}(w)}{(1-H^{t}(v))^{2}}H_{k}^{t}(\mathop{}\!\mathrm{d}u)H_{k}^{0,t}(\mathop{}\!\mathrm{d}v)H_{k}^{1,t}(\mathop{}\!\mathrm{d}w)$
$+2\int\int I(v<w)\frac{\varphi(w)\gamma_{0}^{t}(w)}{1-H^{t}(v)}H^{0,t}_{k}(\mathop{}\!\mathrm{d}v)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w)$ (A.4)
$-\int\int I(v<w)\frac{\varphi(w)\gamma_{0}^{t}(w)}{1-H^{t}(v)}H^{0,t}(\mathop{}\!\mathrm{d}v)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w)+r_{k,n}^{1},$
where
$r_{k,n}^{1}=\int\int\varphi(w)\gamma_{0}^{t}(w)I(z<w)\frac{(H_{k}^{t}(z)-H^{t}(z))^{2}}{(1-H^{t}(z))^{2}(1-H^{t}_{k}(z))}H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w).$ (A.5)
As a consequence of Temporary Condition 1, all denominators appearing in the proof are bounded away from zero. Hereinafter, integrals without are understood to be from 1 to $\infty$.∞ .
Lemma A.2.
Under the above conditions we have that the following term is $\overline{o}_{P}(k^{-1/2})$:
$\int\int I(v<w)\frac{\varphi(w)\gamma_{0}^{t}(w)}{1-H^{t}(v)}[(H^{0,t}_{k}(\mathop{}\!\mathrm{d}v)-H^{0,t}(\mathop{}\!\mathrm{d}v))(H^{1,t}_{k}(\mathop{}\!\mathrm{d}w)-H^{1,t}(\mathop{}\!\mathrm{d}w))].$
Proof.
Let us prove the analogue of Lemma 2.3, [38], in our setting. The proof of the corresponding analogue of Lemma 2.2, [38], is similar. So, consider the second term in the rhs of (A.4)
$V_{k,n}$ $:=\int\int I(v<w)\frac{\varphi(w)\gamma_{0}^{t}(w)}{1-H^{t}(v)}H^{0,t}_{k}(\mathop{}\!\mathrm{d}v)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w)$
$=\frac{1}{k^{2}}\sum_{i=1}^{k}\sum_{j=1}^{k}\frac{I(V_{i}^{\ast}<V_{j}^{\ast})(1-\delta_{i}^{\ast})\delta_{j}^{\ast}\varphi(V_{j}^{\ast})\gamma_{0}^{t}(V_{j}^{\ast})}{1-H^{t}(V_{i}^{\ast})}.$
We now rely on several results from classical U-statistics, specifically, Lemma 5.2.1A and Theorem 5.3.2 in [35]. First, note that $V_{k,n}$ is in fact a V-statistic (see for definition, e.g., Section 5.1.2, [35]). Next, define
$U_{k,n}:$ $=\frac{k}{k-1}V_{k,n}$
$=\frac{1}{k(k-1)}\sum_{i=1}^{k}\sum_{j\neq i}\frac{I(V_{i}^{\ast}<V_{j}^{\ast})(1-\delta_{i}^{\ast})\delta_{j}^{\ast}\varphi(V_{j}^{\ast})\gamma_{0}^{t}(V_{j}^{\ast})}{1-H^{t}(V_{i}^{\ast})}$
$=\frac{1}{k(k-1)}\sum_{i=2}^{k}\sum_{j=1}^{i-1}\left(\frac{I(V_{i}^{\ast}<V_{j}^{\ast})(1-\delta_{i}^{\ast})\delta_{j}^{\ast}\varphi(V_{j}^{\ast})\gamma_{0}^{t}(V_{j}^{\ast})}{1-H^{t}(V_{i}^{\ast})}\right.$
$\qquad\qquad\qquad\qquad+\left.\frac{I(V_{j}^{\ast}<V_{i}^{\ast})(1-\delta_{j}^{\ast})\delta_{i}^{\ast}\varphi(V_{i}^{\ast})\gamma_{0}^{t}(V_{i}^{\ast})}{1-H^{t}(V_{j}^{\ast})}\right)$
$=\frac{1}{k(k-1)}\sum_{c}\left(h_{0}\Big{(}\binom{V_{i}^{\ast}}{\delta_{i}^{\ast}},\binom{V_{j}^{\ast}}{\delta^{\ast}_{j}}\Big{)}+h_{0}\Big{(}\binom{V_{j}^{\ast}}{\delta^{\ast}_{j}},\binom{V_{i}^{\ast}}{\delta_{i}^{\ast}}\Big{)}\right)$
$=\frac{2}{k(k-1)}\sum_{c}h\Big{(}\binom{V_{i}^{\ast}}{\delta_{i}^{\ast}},\binom{V_{j}^{\ast}}{\delta^{\ast}_{j}}\Big{)},$
where $\sum_{c}$ denotes summation over the $\binom{k}{2}$ combinations of 2 distinct elements $\{i_{1},i_{2}\}$ from $\{1,\ldots,k\},$
$h_{0}(\textbf{x},\textbf{y})=\frac{I(x_{1}<y_{1})(1-x_{2})y_{2}\varphi(y_{1})\gamma_{0}^{t}(y_{1})}{1-H^{t}(x_{1})}$
with $\textbf{x},\textbf{y}\in\amsmathbb{R}^{2},$ and $h(\textbf{x},\textbf{y})=\big{[}h_{0}(\textbf{x},\textbf{y})+h_{0}(\textbf{y},\textbf{x})\big{]}/2.$ Clearly, the above expression casts $U_{k,n}$ as a U-statistic, since the kernel $h$ is symmetric. Denote
$\theta^{t}=\amsmathbb{E}h\Big{(}\binom{V_{1}^{\ast}}{\delta_{1}^{\ast}},\binom{V_{2}^{\ast}}{\delta^{\ast}_{2}}\Big{)}.$
Then, according to the definition from Section 5.3.1, [35], the Hajek projection of the U-statistic $U_{k,n}$ is
$\widehat{U}_{k,n}=\sum_{i=1}^{k}\amsmathbb{E}\left[U_{k,n}\left|\binom{V_{i}^{\ast}}{\delta^{\ast}_{i}}\right.\right]-(k-1)\theta^{t}=\frac{2}{k}\sum_{i=1}^{k}h_{1}\Big{(}\binom{V_{i}^{\ast}}{\delta^{\ast}_{i}}\Big{)}-\theta^{t},$
where
$h_{1}(\textbf{x})=\amsmathbb{E}\,h\Big{(}\textbf{x},\binom{V_{1}^{\ast}}{\delta^{\ast}_{1}}\Big{)}.$
Similarly to Section 5.3.2, [35], we have that $U_{k,n}-\widehat{U}_{k,n}$ is also a U-statistic,
$U_{k,n}-\widehat{U}_{k,n}=\frac{1}{\binom{k}{2}}\sum_{c}H\Big{(}\binom{V_{i}^{\ast}}{\delta_{i}^{\ast}},\binom{V_{j}^{\ast}}{\delta^{\ast}_{j}}\Big{)}$
but now with the symmetric kernel
$H(\textbf{x},\textbf{y})=h(\textbf{x},\textbf{y})-h_{1}(\textbf{x})-h_{1}(\textbf{y})+\theta^{t}.$
Following the notation $\zeta_{c}$ from Section 5.2.1, [35], we define in our case for $c=0,1,2$,
$\zeta_{0}=0,\>\>\zeta_{1}={\rm var}\,h_{1}\Big{(}\binom{V_{1}^{\ast}}{\delta^{\ast}_{1}}\Big{)}\>\>\text{ and }\>\>\zeta_{2}={\rm var}\,h\Big{(}\binom{V_{1}^{\ast}}{\delta_{1}^{\ast}},\binom{V_{2}^{\ast}}{\delta^{\ast}_{2}}\Big{)},$ (A.6)
and note that $\zeta_{0}\leq\zeta_{1}\leq\zeta_{2}.$ Define also $\zeta^{H}_{c},$ $c=0,1,2,$ by replacing the kernel $h$ with $H$ in (A.6) and note that $\zeta^{H}_{1}=0,$ see the argument before Theorem 5.3.2 in [35]. Based on (*) in Lemma 5.2.1, [35], we derive
${\rm var}\,U_{k,n}=\frac{4(k-2)}{k(k-1)}\zeta_{1}+\frac{2}{k(k-1)}\zeta_{2},\quad{\rm var}\,\big{(}U_{k,n}-\widehat{U}_{k,n}\big{)}=\frac{2}{k(k-1)}\zeta^{H}_{2}.$ (A.7)
Moreover, by the Cauchy-Schwarz inequality
$\zeta_{2}^{H}\leq 6\zeta_{1}+3\zeta_{2}.$
Contrary to Lemma 2.3 in [38], in our setting ${\rm var}\,U_{k,n}$ and ${\rm var}\,\big{(}U_{k,n}-\widehat{U}_{k,n}\big{)}$ might depend on $t$ asymptotically as $t\to\infty,$ and thus an appeal to (A.7) is not enough to prove $U_{k,n}-\widehat{U}_{k,n}=O(k^{-1})$ in probability when transitioning from the conditional setting given $Z_{n-k,n}=t$ to the unconditional setting.
We now argue that ${\rm var}\,U_{k,n}$ and ${\rm var}\,\big{(}U_{k,n}-\widehat{U}_{k,n}\big{)}$ do not depend on $t$ asymptotically as $t\to\infty$ and $\zeta_{1},$ $\zeta_{2}$ can be bounded from above uniformly in large $t$, and thus, the arguments of Lemmas 2.2 and 2.3, [38], may be adapted to our setting. Indeed, for $\theta^{t},$ we get by (2.8), (2.11), Potter’s bounds, and the dominated convergence theorem
$\theta^{t}$ $=\int_{1}^{\infty}\int_{1}^{\infty}\frac{I(v<w)\varphi(w)\gamma_{0}^{t}(w)}{1-H^{t}(v)}H^{0,t}(\mathop{}\!\mathrm{d}v)H^{1,t}(\mathop{}\!\mathrm{d}w)$
$=\int_{1}^{\infty}\int_{1}^{w}\varphi(w)\frac{1}{(1-G^{t}(v))}G^{t}(\mathop{}\!\mathrm{d}v)F^{t}(\mathop{}\!\mathrm{d}w)$
$\to\frac{1}{\gamma_{G}\gamma_{F}}\int_{1}^{\infty}\varphi(w)\log(w)w^{-1/\gamma_{F}-1}\mathop{}\!\mathrm{d}w=:\theta,$ (A.8)
Let us explain the latter convergence in more detail since we use the same argument repeatedly in the proof. The integral in the second line of (A.8) can be represented as a mathematical expectation of some random variable depending on $t.$ To apply the dominated convergence theorem, we should show that this random variable is bounded uniformly in large $t$ by some random variable with finite expectation. By Potter’s bounds, eventual continuity of $G$ and continuity of $\bar{\varphi},$ for some $\delta\in(0,1)$ and large $t$
$|\varphi(w)|\int_{1}^{w}\frac{1}{(1-G^{t}(v))}G^{t}(\mathop{}\!\mathrm{d}v)\leq-\bar{\varphi}(w)\log(1-G^{t}(w))$
$\leq\bar{\varphi}(w)\big{(}\log w\,(1/\gamma_{G}+\delta)-\log(1-\delta)\big{)}$ (A.9)
$\leq A\,I(w\in[1,1+\delta])+B\,I(w>1+\delta)\bar{\varphi}(w)\log w,$
for some positive constants $A$ and $B$. Recall that $U_{F}$ denotes the tail quantile function of $F,$ i.e. $U_{F}(x):=\inf\{y:F(y)\geq 1-1/x\}.$ Then one can show that
$U_{F^{t}}(x)=\frac{U_{F}(x/(1-F(t)))}{U_{F}(1/(1-F(t)))}.$ (A.10)
By Potter’s bounds, we get for some $\bar{\delta},\varepsilon>0,$ $T_{0}>1$ and all $t>T_{0}$
$f_{\ast}(x):=(1-\bar{\delta})x^{\gamma_{F}-\varepsilon}\leq\frac{U_{F}(x/(1-F(t)))}{U_{F}(1/(1-F(t)))}\leq(1+\bar{\delta})x^{\gamma_{F}+\varepsilon}=f^{\ast}(x).$ (A.11)
Note also that
$\int_{1}^{\infty}\bar{\varphi}(w)\log w\,F^{t}(\mathop{}\!\mathrm{d}w)=\amsmathbb{E}\big{[}\bar{\varphi}(U_{F^{t}}(\xi))\log(U_{F^{t}}(\xi))\big{]},$
where $\xi$ denotes a Pareto$(1)$ random variable. Assume that $\bar{\varphi}(x)$ is monotone for $x>T_{1}\geq 1.$ Moreover, assume without loss of generality that $\bar{\varphi}(x)$ is non-decreasing on this set. Denote $T_{2}=\max(T_{0},T_{1}).$ Then by the two latter identities and properties of $\bar{\varphi}$
$\bar{\varphi}(U_{F^{t}}(w))\log(U_{F^{t}}(w))$$\leq$ $CI(w\in[1,T_{2}])+\bar{\varphi}(U_{F^{t}}(w))\log(U_{F^{t}}(w))I(w>T_{2})$
$\leq$ $CI(w\in[1,T_{2}])+\bar{\varphi}(f^{\ast}(w))\log(f^{\ast}(w))I(w>T_{2})$
for some $C>0,$ where the latter function does not depend on $t$ as required. Finally,
$\amsmathbb{E}\big{[}\bar{\varphi}(f^{\ast}(\xi))\log(f^{\ast}(\xi))I(\xi>T_{2})\big{]}$ $=$ $\int_{T_{2}}^{\infty}\bar{\varphi}(f^{\ast}(x))\log(f^{\ast}(x))\frac{1}{x^{2}}\mathop{}\!\mathrm{d}x$
$\leq$ $C_{1}\int_{T_{3}}^{\infty}\bar{\varphi}(x)\log x\,x^{-1/(\gamma_{F}+\varepsilon)-1}\mathop{}\!\mathrm{d}x,$
for some $C_{1}>0$ and $T_{3}>1,$ and the last integral is finite by continuity of $\bar{\varphi}$ and Temporary Condition 1. Summarizing the above, for all large $t$ the random variable
$\eta_{t}=|\varphi(U_{F^{t}}(\xi))|\int_{1}^{U_{F^{t}}(\xi)}\frac{1}{(1-G^{t}(v))}G^{t}(\mathop{}\!\mathrm{d}v)$
is bounded by some integrable random variable, therefore the conditions of the dominated convergence theorem hold, and the latter implies (A.8).
In a similar fashion, applying the dominated convergence theorem and using additionally Potter’s bounds for $H^{t},$ we get the uniform upper bound for $\zeta_{2}$ and thus,
$\zeta_{2}$ $=$ $\int_{1}^{\infty}\int_{1}^{\infty}\frac{I(v<w)\varphi^{2}(w)(\gamma_{0}^{t}(w))^{2}}{(1-H^{t}(v))^{2}}H^{0,t}(\mathop{}\!\mathrm{d}v)H^{1,t}(\mathop{}\!\mathrm{d}w)-(\theta^{t})^{2}$
$=$ $\int_{1}^{\infty}\int_{1}^{w}\varphi^{2}(w)\frac{1-F^{t}(v)}{(1-H^{t}(v))^{2}(1-G^{t}(w))}G^{t}(\mathop{}\!\mathrm{d}v)F^{t}(\mathop{}\!\mathrm{d}w)-(\theta^{t})^{2}$
$\to$ $\frac{\gamma_{H}}{\gamma_{G}\gamma_{F}}\int_{1}^{\infty}\varphi^{2}(w)w^{1/\gamma_{H}-1}\mathop{}\!\mathrm{d}w-\theta^{2}.$
The analysis of $\zeta_{1}$ is similar, thus we omit it.
To conclude, by a final application of Potter’s bounds, we have established that the variance of $U_{k,n}-\widehat{U}_{k,n}$ is $\overline{O}_{P}(k^{-2})$, and thus the proof of Lemma 2.3, [38] (and, consequently, Lemma 2.2 ibid., see Lemma A.3 below), can be reused in our setting. In other words, as $k,\>n/k\to\infty$, the random sequence $U_{k,n}-\widehat{U}_{k,n}$ is $\overline{o}_{P}(k^{-1/2})$. ∎
We may prove an analogous result for the first term in the rhs of (A.4).
Lemma A.3.
Under the above conditions we have that the following term is $\overline{o}_{P}(k^{-1/2})$:
$\int\int\int\frac{I(v<u,v<w)\varphi(w)\gamma_{0}^{t}(w)}{(1-H^{t}(v))^{2}}\cdot[H_{k}^{t}(\mathop{}\!\mathrm{d}u)H_{k}^{0,t}(\mathop{}\!\mathrm{d}v)H_{k}^{1,t}(\mathop{}\!\mathrm{d}w)-H_{k}^{t}(\mathop{}\!\mathrm{d}u)H^{0,t}(\mathop{}\!\mathrm{d}v)H^{1,t}(\mathop{}\!\mathrm{d}w)$
$-H^{t}(\mathop{}\!\mathrm{d}u)H_{k}^{0,t}(\mathop{}\!\mathrm{d}v)H^{1,t}(\mathop{}\!\mathrm{d}w)-H^{t}(\mathop{}\!\mathrm{d}u)H^{0,t}(\mathop{}\!\mathrm{d}v)H_{k}^{1,t}(\mathop{}\!\mathrm{d}w)+2H^{t}(\mathop{}\!\mathrm{d}u)H^{0,t}(\mathop{}\!\mathrm{d}v)H^{1,t}(\mathop{}\!\mathrm{d}w)].$
Proof.
Similar to the previous case. ∎
Thus, similarly to Corollary 2.4, [38], we derive that
$\frac{1}{k}\sum_{i=1}^{k}\varphi(V_{i}^{\ast})\gamma^{t}_{0}(V_{i}^{\ast})\delta^{\ast}_{i}C_{in}^{t}=\frac{1}{k}\sum_{i=1}^{k}\gamma_{1}^{t}(V_{i}^{\ast})(1-\delta^{\ast}_{i})-\frac{1}{k}\sum_{i=1}^{k}\gamma_{2}^{t}(V_{i}^{\ast})+r^{1}_{k,n}+r^{2}_{k,n},$ (A.12)
where $r^{2}_{k,n}=\overline{o}_{P}(k^{-1/2})$.
We turn our attention to the behavior of $r_{k,n}^{1},$ defined by (A.5) (compare with $R_{n1}$ in [38]).
Lemma A.4.
We have that
$|r_{k,n}^{1}|=\overline{O}_{P}(k^{-1}).$ (A.13)
Proof.
First, by Temporary Condition 1
$|r_{k,n}^{1}|$ $\leq\int_{1}^{\infty}|\varphi(w)|\gamma_{0}^{t}(w)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w)\cdot\int_{1}^{\infty}I(z<T)\frac{(H_{k}^{t}(z)-H^{t}(z))^{2}}{(1-H^{t}(z))^{2}(1-H^{t}_{k}(z))}H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)$
$=:I_{k,1}\cdot I_{k,2},$ (A.14)
therefore we can evaluate the asymptotic of $I_{k,1}$ and $I_{k,2}$ separately. First, for every $\varepsilon>0$ by Markov’s inequality and an argument similar to that used for proving (A.8)
${\amsmathbb{P}}(I_{k,1}>\varepsilon)$ $={\amsmathbb{P}}\left(\frac{1}{k}\sum_{i=1}^{k}|\varphi(V_{i}^{\ast})|\gamma_{0}^{t}(V_{i}^{\ast})\delta_{i}^{\ast}>\varepsilon\right)$
$\leq\frac{1}{\varepsilon}\int_{1}^{\infty}|\varphi(w)|F^{t}(\mathop{}\!\mathrm{d}w)\leq\frac{1}{\varepsilon\gamma_{F}}\int_{1}^{\infty}|\bar{\varphi}(w)|w^{-1/\gamma_{F}-1+\delta}\mathop{}\!\mathrm{d}w$
for large $t$ and some $\delta>0.$ Hence we derive that $I_{k,1}=\overline{O}_{P}(1).$
Next, consider $I_{k,2}.$ By (2.12) and Potter’s bounds, we can bound the denominator in $I_{k,2}$ from below by some constant independent of $t$ since $(1-H^{t}(z))/(1-H_{k}^{t}(z))$ is stochastically bounded under the condition $z<T.$ Hence, it is enough to find the asymptotic of the following integral
$I_{k,3}:=\int_{1}^{\infty}I(z<T)(H_{k}^{t}(z)-H^{t}(z))^{2}H^{0,t}_{k}(\mathop{}\!\mathrm{d}z).$
Clearly,
$I_{k,3}\leq\int_{1}^{\infty}(H_{k}^{t}(z)-H^{t}(z))^{2}H^{t}_{k}(\mathop{}\!\mathrm{d}z).$
By the classical goodness-of-fit technique, the latter integral equals
$\int_{0}^{1}(Q_{k}(x)-x)^{2}Q_{k}(\mathop{}\!\mathrm{d}x)=I_{k,4},$
and thus does not depend on $t,$ where $Q_{k}(x)=\frac{1}{k}\sum_{i=1}^{k}I(U_{i}\leq x),$ the empirical cdf of independent $U[0,1]$ random variables $\{U_{i}\}_{i=1}^{k},$ and
$U_{i}=H^{t}(V_{i}^{\ast}),\quad i=1,...,k.$ (A.15)
Finally,
$I_{k,4}=\int_{0}^{1}(Q_{k}(x)-x)^{2}\mathop{}\!\mathrm{d}x+\int_{0}^{1}(Q_{k}(x)-x)^{2}\mathop{}\!\mathrm{d}(Q_{k}(x)-x),$
where the first summand on the rhs is exactly the Cramér-von Mises statistic divided by $k$ and the second summand is asymptotically smaller in probability than the first one, since $\sqrt{k}(Q_{k}(x)-x)$ converges in distribution to a Brownian bridge on $[0,1]$ in a Skorokhod space. Hence, $I_{k,4}$ and therefore $I_{k,2}$ are $\overline{O}_{P}(k^{-1})$ as desired. ∎
Next, consider $S_{k,n}^{1},$ given by
$S_{k,n}^{1}=\frac{1}{k}\sum_{i=1}^{k}\varphi(V_{i}^{\ast})\gamma_{0}^{t}(V_{i}^{\ast})\delta_{i}^{\ast}B_{ik}^{t},$
(compare with $S_{n1}$ from Lemma 2.6 in [38]).
Lemma A.5.
We have that
$S_{k,n}^{1}=\overline{O}_{P}(k^{-1}).$ (A.16)
Proof.
Similarly to the relation (2.4) in [38],
$-\frac{1}{2k}\int_{1}^{V_{i}^{\ast}-}\frac{H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)}{(1-H^{t}_{k}(z))^{2}}\leq B_{ik}^{t}\leq 0.$ (A.17)
The analysis of $S_{k,n}^{1}$ resembles that of $r_{k,n}^{1}.$ We have by (A.14) and (A.17),
$|S_{k,n}^{1}|$$\leq$ $\frac{1}{2k}\int_{1}^{\infty}\int_{1}^{\infty}|\varphi(w)|\gamma_{0}^{t}(w)I(z<w)\frac{1}{(1-H_{k}^{t}(z))^{2}}H_{k}^{0,t}(\mathop{}\!\mathrm{d}z)H_{k}^{1,t}(\mathop{}\!\mathrm{d}w)$
$\leq$ $\frac{1}{2k}I_{k,1}\cdot\int_{1}^{\infty}I(z<T)\frac{1}{(1-H_{k}^{t}(z))^{2}}H_{k}^{0,t}(\mathop{}\!\mathrm{d}z)=:\frac{1}{2k}I_{k,1}\cdot I_{k,5},$
where $I_{k,1}$ was defined in (A.14) and proved to be $\overline{O}_{P}(1).$ Recalling that $Q_{k}$ is the empirical cdf of the $U[0,1]$ random variables defined by (A.15) and does not depend on $t$, we have by Potter’s bound
$I_{k,5}$$\leq$ $\int_{1}^{\infty}I(z<T)\frac{1}{(1-H_{k}^{t}(z))^{2}}H_{k}^{t}(\mathop{}\!\mathrm{d}z)$ (A.18)
$=$ $\int_{0}^{1}I(x<H^{t}(T))\frac{1}{(1-Q_{k}(x))^{2}}Q_{k}(\mathop{}\!\mathrm{d}x)$
$\leq$ $\int_{0}^{1}I(x<1-(1-\delta)T^{-1/\gamma_{H}-\delta})\frac{1}{(1-Q_{k}(x))^{2}}Q_{k}(\mathop{}\!\mathrm{d}x)$
for large $t$ and some $\delta>0,$ which implies the result. ∎
Finally, consider $S_{k,n}^{2},$ given by
$S_{k,n}^{2}=\frac{1}{2k}\sum_{i=1}^{k}|\varphi(V_{i}^{\ast})|\delta^{\ast}_{i}e^{\Delta_{i}^{t}}\left[B_{ik}^{t}+C_{ik}^{t}\right]^{2},$ (A.19)
(compare with $S_{n2}$ from Lemma 2.7 [38]).
Lemma A.6.
We have that
$S_{k,n}^{2}=\overline{O}_{P}(k^{-1}).$ (A.20)
Proof.
In fact, (A.17) and (A.18) imply $B_{ik}^{t}=\overline{O}_{P}(k^{-1})$. Let us show that $C_{ik}^{t}=\overline{O}_{P}(k^{-1/2}).$ From (A.3) and (A.15) it follows
$C_{ik}^{t}=\int_{0}^{U_{i}-}\frac{1}{1-Q_{k}(x)}Q^{0}_{k}(\mathop{}\!\mathrm{d}x)-\int_{0}^{U_{i}-}\frac{1}{1-x}Q^{0}(\mathop{}\!\mathrm{d}x),$
where
$Q^{0}_{k}(x)=\frac{1}{k}\sum_{i=1}^{k}I(U_{i}\leq x,\delta_{i}^{\ast}=0)\quad\text{ and }\quad Q^{0}(x)={\amsmathbb{P}}(U_{1}\leq x,\delta_{1}^{\ast}=0).$
We have,
$C_{ik}^{t}$ $\leq\left|\int_{0}^{U_{i}-}\left(\frac{1}{1-Q_{k}(x)}-\frac{1}{1-x}\right)Q^{0}_{k}(\mathop{}\!\mathrm{d}x)\right|+\left|\int_{0}^{U_{i}-}\frac{1}{1-x}Q^{0}_{k}(\mathop{}\!\mathrm{d}x)-\int_{0}^{U_{i}-}\frac{1}{1-x}Q^{0}(\mathop{}\!\mathrm{d}x)\right|$
$=:C_{ik}^{t1}+C_{ik}^{t2}.$
Note that we can consider only $C_{ik}^{t}$ with $U_{i}\leq H^{t}(T)$ in (A.19) due to Temporary Condition 1, where $H^{t}(T)\leq T_{0}$ for large $t$ and some $T_{0}\in(0,1)$ by (A.18). Thus, by Kolmogorov’s theorem and the fact that $(1-x)/(1-Q_{k}(x))$ is stochastically bounded from above on $x<U_{k,k}$ (see p. 415, [36]), for large $k$
$\sup_{x\in[0,U_{i})}\left|\frac{1}{1-Q_{k}(x)}-\frac{1}{1-x}\right|\leq\frac{\sup_{x\in[0,T_{0}]}|Q_{k}(x)-x|}{(1-Q_{k}(T_{0}))(1-T_{0})}=\overline{O}_{P}(k^{-1/2})$
since neither $Q_{k}(x)$ nor $x$ depend on $t.$ This implies $C_{ik}^{t1}=\overline{O}_{P}(k^{-1/2})$.
Let us analyse $C_{ik}^{t2}.$ The main difficulty is that $Q^{0}_{k}$ and $Q^{0}$ depend on $t$ via $\{\delta^{\ast}_{i}\}.$ First, let us equivalently represent the integrals in $C_{ik}^{t2}$ as integrals with respect to proper cdfs (and not only by sub-distribution functions). Consider the auxiliary random variables $\bar{U}_{i},i=1,...,k,$ where
$\bar{U}_{i}=U_{i}(1-\delta_{i}^{\ast})+\delta_{i}^{\ast}(1-\varepsilon_{i}),$
and $\{\varepsilon_{i}\}_{i=1}^{k}$ are independent random variables, uniformly distributed on $[0,\bar{\varepsilon}]$ with $\bar{\varepsilon}<1-T_{0}$ and independent of $\{U_{i}\}_{i=1}^{k}.$ Denote the cdf and empirical cdf of $\{\bar{U}_{i}\}_{i=1}^{k}$ by $\bar{Q}(x)$ and $\bar{Q}_{k}(x),$ respectively. Then, for all $x\in[0,T_{0}]$
$Q^{0}(x)=\bar{Q}(x)\quad\text{ and }\quad Q^{0}_{k}(x)=\bar{Q}_{k}(x).$
Next, let a non-increasing step function $f_{k}:[0,1)\to[1,\infty)$ and a partition $\{\Delta_{j}\}_{j=1}^{\infty}$ of the interval $[0,1)$ be such that
$f_{k}(x)=1+\frac{j-1}{k}\quad\text{and}\quad 0\leq\frac{1}{1-x}-f_{k}(x)<\frac{1}{k}\quad\text{for }x\in\Delta_{j},\;j\in\amsmathbb{N}.$
Clearly, $f_{k}$ and $\{\Delta_{j}\}_{j=1}^{\infty}$ are defined uniquely by the latter. Then for every $j\in\amsmathbb{N}$ such that $\Delta_{j}\cap[0,T_{0}]\neq\emptyset$
$\left|\int_{\Delta_{j}}\frac{1}{1-x}\big{(}\bar{Q}_{k}(\mathop{}\!\mathrm{d}x)-\bar{Q}(\mathop{}\!\mathrm{d}x)\big{)}\right|$$\leq$ $\left|\int_{\Delta_{j}}f_{k}(x)\big{(}\bar{Q}_{k}(\mathop{}\!\mathrm{d}x)-\bar{Q}(\mathop{}\!\mathrm{d}x)\big{)}\right|+\frac{1}{k}\left|\int_{\Delta_{j}}\big{(}\bar{Q}_{k}(\mathop{}\!\mathrm{d}x)-\bar{Q}(\mathop{}\!\mathrm{d}x)\big{)}\right|$ (A.21)
$\leq$ $2(C+k^{-1})\sup_{x\in[0,1]}\big{|}\bar{Q}_{k}(x)-\bar{Q}(x)\big{|},$
where $C=kf_{k}(T_{0}).$ Denote $x_{j-1}$ and $x_{j},$ the endpoints of $\Delta_{j},$ $j\in\amsmathbb{N},$ and notice that $x_{0}=0.$ By (A.21) we have
$\int_{0}^{U_{i}-}\frac{1}{1-x}\big{(}\bar{Q}_{k}(\mathop{}\!\mathrm{d}x)-\bar{Q}(\mathop{}\!\mathrm{d}x)\big{)}$ $=$ $\sum_{j=1}^{kf_{k}(U_{i}-)}\Big{(}1+\frac{j-1}{k}\Big{)}\big{(}\bar{Q}_{k}(x_{j})-\bar{Q}_{k}(x_{j-1})-\bar{Q}(x_{j})+\bar{Q}(x_{j-1})\big{)}$
$+kf_{k}(U_{i}-)\cdot O_{P}(1/k)\sup_{x\in[0,T_{0}]}\big{|}\bar{Q}_{k}(x)-\bar{Q}(x)\big{|}$
$=$ $f_{k}(U_{i}-)\big{(}\bar{Q}_{k}(x_{m})-\bar{Q}(x_{m})\big{)}+O_{P}(1)\sup_{x\in[0,T_{0}]}\big{|}\bar{Q}_{k}(x)-\bar{Q}(x)\big{|}$
$+\frac{1}{k}\sum_{j=1}^{kf_{k}(U_{i}-)}\big{(}\bar{Q}_{k}(x_{j})-\bar{Q}_{k}(x_{j-1})\big{)}$
$=$ $O_{P}(1)\sup_{x\in[0,T_{0}]}\big{|}\bar{Q}_{k}(x)-\bar{Q}(x)\big{|},$
where $m$ denotes $kf_{k}(U_{i}-),$ the final relation holds since $f_{k}(U_{i}-)\leq f_{k}(T_{0})<\infty,$ and $O_{P}(1)$ in the last line does not depend on $t.$ By another application of Kolmogorov’s theorem,
$\sup_{x\in[0,T_{0}]}|\bar{Q}_{k}(x)-\bar{Q}(x)|=\overline{O}_{P}(k^{-1/2}),$
since the latter supremum does not depend on $t$ in distribution. This proves $C_{ik}^{t}=\overline{O}_{P}(k^{-1/2})$.
Next, it follows from the asymptotics of $B_{ik}^{t}$ and $C_{ik}^{t}$ that $e^{\Delta_{i}}=\overline{O}_{P}(1)$. Finally,
$\frac{1}{k}\sum_{i=1}^{k}|\varphi(V_{i}^{\ast})|\delta^{\ast}_{i}\leq\frac{1}{k}\sum_{i=1}^{k}|\varphi(V_{i}^{\ast})|\gamma_{0}^{t}(V_{i}^{\ast})\delta^{\ast}_{i}=\overline{O}_{P}(1).$
Summarizing the above, we have found that $S^{2}_{k,n}=\overline{O}_{P}(k^{-1})$, as desired. ∎
Piecing all the above asymptotics together, that is appealing to (A), (A.12), (A.13), (A.16), and (A.20) completes the proof of Theorem 3.1. ∎
Proof of Proposition 3.2.
By eventual continuity of $F$ and $G,$
$\int_{1}^{\infty}\varphi^{2}(x)(\gamma^{t}_{0}(x))^{2}H^{1,t}(\mathop{}\!\mathrm{d}x)=\int_{1}^{\infty}\varphi^{2}(x)\frac{F^{t}(\mathop{}\!\mathrm{d}x)}{1-G^{t}(x)}.$
By regular variation property of $F$ and $G,$ and the argument similar to that used to prove (A.8), under Condition 1 we derive that
$\int_{1}^{\infty}\varphi^{2}(x)(\gamma^{t}_{0}(x))^{2}H^{1,t}(\mathop{}\!\mathrm{d}x)\leq C\int_{1}^{\infty}\bar{\varphi}^{2}(x)x^{1/\gamma_{G}-1/\gamma_{F}-1+\varepsilon}\mathop{}\!\mathrm{d}x<\infty$ (A.22)
for some $C>0$ and large $t,$ so (3.6) is satisfied. Next, it is easy to show that $C^{t}(x)$ is regularly varying with index $1/\gamma_{H},$ thus by the same argument for some $C>0$ and large $t$
$\int_{1}^{\infty}|\varphi(x)|\sqrt{C^{t}(x)}F^{t}(\mathop{}\!\mathrm{d}x)$$\leq$ $C\int_{1}^{\infty}\bar{\varphi}(x)x^{(2\gamma_{G})^{-1}-(2\gamma_{F})^{-1}-1+\varepsilon/2}\mathop{}\!\mathrm{d}x$
$\leq$ $C\left(\int_{1}^{\infty}\bar{\varphi}^{2}(x)x^{1/\gamma_{G}-1/\gamma_{F}-2+\varepsilon}\mathop{}\!\mathrm{d}x\right)^{1/2}<\infty,$
and so (3.7) is satisfied as well. ∎
Proof of Theorem 3.3.
The main steps of the proof use the last steps of the proof of Theorem 1.1 in [38]. First, for a given $\epsilon>0$ let us choose a function $\tilde{\varphi}$ vanishing outside of $[1,T],$ for some $T<\infty,$ and satisfying Condition 1, such that
$\int_{1}^{\infty}(\overline{(\varphi-\tilde{\varphi})}(x))^{2}x^{\alpha(\varepsilon)}\mathop{}\!\mathrm{d}x\leq\epsilon,$ (A.24)
where $\overline{\varphi-\tilde{\varphi}}$ denotes an envelope of the function $\varphi-\tilde{\varphi},$ which is possible under Condition 1. Indeed, it suffices to choose $\tilde{\varphi}(x)=\varphi(x)I(x\leq T),$ $x\geq 1.$ Note that by (A.24) it is possible to find such $\epsilon_{1}>0,$ vanishing together with $\epsilon$, such that
$\int_{1}^{\infty}(\varphi-\tilde{\varphi})^{2}(\gamma^{t}_{0})^{2}\mathop{}\!\mathrm{d}H^{1,t}\leq\epsilon_{1}$ (A.25)
and
$\int_{1}^{\infty}|\varphi-\tilde{\varphi}|\sqrt{C^{t}}\mathop{}\!\mathrm{d}F^{t}\leq\epsilon_{1}$ (A.26)
for large $t$ similarly to (A.22) and (A), respectively.
Set $\varphi_{1}=\varphi-\tilde{\varphi}.$ We prove that
$k^{1/2}\left(\int_{1}^{\infty}\varphi_{1}\mathop{}\!\mathrm{d}\amsmathbb{F}^{t}_{k,n}-\int_{1}^{\infty}\varphi_{1}\mathop{}\!\mathrm{d}F^{t}\right)=\overline{O}_{P}(\epsilon_{1}^{1/2}),$ (A.27)
which together with Theorem 3.1 and the bound
$\int_{1}^{\infty}|\varphi_{1}|\mathop{}\!\mathrm{d}F^{t}<\epsilon_{1}$
following from (A.26) implies the conclusion of Theorem 3.3. Similarly to the first formula on p.436 in [38], we get
$k^{1/2}\int_{1}^{\infty}\varphi_{1}\,\mathop{}\!\mathrm{d}(\amsmathbb{F}^{t}_{k,n}-F^{t})$ $=k^{-1/2}\sum_{i=1}^{k}\left[\varphi_{1}(V_{i}^{\ast})\delta_{i}^{\ast}\gamma_{0}^{t}(V_{i}^{\ast})-\int_{1}^{\infty}\varphi_{1}\mathop{}\!\mathrm{d}F^{t}\right]$
$\quad+k^{-1/2}\sum_{i=1}^{k}\varphi_{1}(V_{i}^{\ast})\delta_{i}^{\ast}\gamma_{0}^{t}(V_{i}^{\ast})\big{[}\exp(B_{ik}^{t}+C_{ik}^{t})-1\big{]}.$
By (A.25), the variance of the first summand in the right-hand side of the latter relation is not more than $2\epsilon_{1},$ thus by Chebyshev’s inequality the first summand is $\overline{O}_{P}(\epsilon_{1}^{1/2}).$ By the trivial bound $|\exp(x)-1|\leq|x|\exp(|x|),$ the second summand is bounded in absolute value by
$k^{-1/2}\sum_{i=1}^{k}|\varphi_{1}(V_{i}^{\ast})|\delta_{i}^{\ast}\gamma_{0}^{t}(V_{i}^{\ast})(|B_{ik}^{t}|+|C_{ik}^{t}|)\exp\big{[}|B_{ik}^{t}|+|C_{ik}^{t}|\big{]}.$
By (A.17),
$|B_{ik}^{t}|\leq\frac{1}{2k}\int_{1}^{V_{i}^{\ast}-}\frac{H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)}{(1-H^{t}_{k}(z))^{2}}\leq\frac{1}{2k}\int_{1}^{V_{i}^{\ast}-}\frac{H^{t}_{k}(\mathop{}\!\mathrm{d}z)}{(1-H^{t}_{k}(z))^{2}}\leq\frac{1}{k(1-H^{t}_{k}(V^{\ast}_{k,k}-))}\leq 1.$
Now let us prove that
$\sup_{1\leq i\leq k}|C_{ik}^{t}|=\overline{O}_{P}(1)$ (A.28)
as well. Indeed,
$\sup_{1\leq i\leq k}|C_{ik}^{t}|\leq\sup_{z\in[1,V^{\ast}_{k,k})}\left|\int_{1}^{z}\frac{H^{0,t}_{k}(\mathop{}\!\mathrm{d}x)}{1-H^{t}_{k}(x)}-\int_{1}^{z}\frac{H^{0,t}(\mathop{}\!\mathrm{d}x)}{1-H^{t}(x)}\right|.$
Theorem 2.1 in [45] yields an upper bound on the exceedance probability of the supremum on the right-hand side of the latter relation over some level $a$ that depends on $a$ alone (concretely, the bound can be taken to be $358\cdot a^{-2/3}$) and does not depend on distributional characteristics, which implies (A.28).
Thus, since we proved $\sup_{1\leq i\leq k}|B_{ik}^{t}|+|C_{ik}^{t}|=\overline{O}_{P}(1),$ it remains to bound the following term from above
$k^{-1/2}\sum_{i=1}^{k}|\varphi_{1}(V_{i}^{\ast})|\delta_{i}^{\ast}\gamma_{0}^{t}(V_{i}^{\ast})(|B_{ik}^{t}|+|C_{ik}^{t}|),$
which can be done through the same reasoning as in the corresponding part of the proof of Theorem 1.1 in [38] replacing the assumption (2.10) therein with (A.26), and obtaining a bound which is equal to a constant times $\epsilon_{1}$, which in particular is $\overline{O}_{P}(\epsilon_{1})=\overline{o}_{P}(\epsilon_{1}^{1/2})$. ∎
Proof of Theorem 3.4.
We use the proof technique from Theorem 2.1 in [34]. By Theorem 3.3, we have
$S_{k,n}(\varphi)=\frac{1}{k}\sum_{j=1}^{k}W_{j}+r_{k,n},$ (A.29)
where $\{W_{j}\}_{j=1}^{k}$ are i.i.d. given $Z_{n-k,n}=t,$ and $\sqrt{k}r_{k,n}$ is $\overline{o}_{P}(1)$, and thus the last term does not affect the asymptotic of the EKM integral. Note that by (A.27), we can consider $\tilde{\varphi}=\varphi I(\cdot\leq T_{0})$ instead of $\varphi$ in (A.29) for some large $T_{0},$ since $\varphi_{1}=\varphi-\tilde{\varphi}$ does not affect the asymptotic of $S_{k,n}(\varphi)$ by Theorem 3.3.
Denote
$W^{\ast}=\varphi(V^{\ast})\gamma_{0}^{t}(V^{\ast})\delta^{\ast}+\gamma_{1}^{t}(V^{\ast})(1-\delta^{\ast})-\gamma_{2}^{t}(V^{\ast})$ (A.30)
with $(V^{\ast},\delta^{\ast})\stackrel{{\scriptstyle d}}{{=}}(V_{1}^{\ast},\delta_{1}^{\ast})$ and consider the characteristic function of the normalized sum in (A.29) given $Z_{n-k,n}=t.$ By (3.4), we have,
$b_{k}^{t}(s)$ ::=:= $\amsmathbb{E}\left[\left.\exp\Big{(}\frac{is}{k}\sum_{j=1}^{k}W_{j}\Big{)}\right|Z_{n-k,n}=t\right]$
$=$ $\left(\amsmathbb{E}\left[\left.\exp\Big{(}\frac{is}{k}W_{1}\Big{)}\right|Z_{n-k,n}=t\right]\right)^{k}=\left(\amsmathbb{E}\left[\exp\Big{(}\frac{is}{k}W^{\ast}\Big{)}\right]\right)^{k}$
By properties of characteristic functions (more precisely, here we use Lemma A.1, [34]),
$\amsmathbb{E}\big{[}\exp(iz\,W^{\ast})\big{]}=1+iz\amsmathbb{E}W^{\ast}+\eta(z,t),$ (A.31)
with $\eta(z,t)\leq 2|z|^{1+\delta}\amsmathbb{E}|W^{\ast}|^{1+\delta}$ and some $\delta>0$ which we may specify at a later stage. As mentioned above, replacing $\varphi$ with $\tilde{\varphi},$ does not change the asymptotic behaviour of $S_{k,n}(\varphi),$ however, the boundedness of $\amsmathbb{E}|W^{\ast}|^{1+\delta}$ for large $t$ is trivial after such a replacement. Moreover, applying the argument used to prove (A.8), we derive that $\amsmathbb{E}|W^{\ast}|^{1+\delta}$ is bounded uniformly in $t\geq T.$ Thus, for $t$ larger than some $T>1,$ $\eta(z,t)$ is bounded by $m|z|^{1+\delta}$ for some positive constant $m$ independent of $t.$ But then for every $s\in\amsmathbb{R}$
$b_{k}^{t}(s)\to\exp(is\,\amsmathbb{E}W^{\ast})\text{ as }k\to\infty\text{ uniformly in }t\geq T.$ (A.32)
Now, it is straightforward to see that $\amsmathbb{E}\big{[}\gamma_{1}^{t}(V^{\ast})(1-\delta^{\ast})\big{]}=\amsmathbb{E}\gamma^{t}_{2}(V^{\ast}),$ and thus by Condition 1, the argument used for proving (A.8), and dominated convergence theorem
$\amsmathbb{E}W^{\ast}=\amsmathbb{E}\big{[}\varphi(V^{\ast})\gamma_{0}^{t}(V^{\ast})\delta^{\ast}\big{]}=\int_{1}^{\infty}\varphi(x)F^{t}(\mathop{}\!\mathrm{d}x)\to\int_{1}^{\infty}\varphi(x)F^{\circ}(\mathop{}\!\mathrm{d}x),\quad t\to\infty.$ (A.33)
Finally, we can write
$\amsmathbb{E}\left[\exp\Big{(}\frac{is}{k}\sum_{j=1}^{k}W_{j}\Big{)}\right]=\amsmathbb{E}[b_{k}^{Z_{n-k,n}}(s)I(Z_{n-k,n}<T)]+\amsmathbb{E}[b_{k}^{Z_{n-k,n}}(s)I(Z_{n-k,n}\geq T)],$
where the first term vanishes since $Z_{n-k,n}\stackrel{{\scriptstyle P}}{{\to}}\infty$ and $|b_{k}^{t}(s)|\leq 1,$ and the second term converges to $\exp(is\,S_{\circ}(\varphi))$ by (A.32) and (A.33). This completes the proof. ∎
Proof of Theorem 3.6.
The ideas of the proofs of Theorems 3.4 and 3.6 are similar. Since $r_{k,n}$ in the decomposition (A.29) is $\overline{o}_{P}(k^{-1/2})$ by Theorem 3.3, it does not play a role in the asymptotic behaviour of $\sqrt{k}S_{k,n}(\varphi).$ Denote for brevity
$S_{t}(\varphi)=\int_{1}^{\infty}\varphi\mathop{}\!\mathrm{d}F^{t},$
and consider the characteristic function of the left-hand side of (3.9) minus $\sqrt{k}r_{k,n}=\overline{o}_{P}(1)$ given $Z_{n-k,n}=t.$ We have
$c_{k}^{t}(s)$ $:=\amsmathbb{E}\left[\left.\exp\Big{(}is\,\sqrt{k}\Big{[}\frac{1}{k}\sum_{i=1}^{k}W_{i}-\int\varphi\,\mathop{}\!\mathrm{d}F^{Z_{n-k,n}}\Big{]}\Big{)}\right|Z_{n-k,n}=t\right]$ (A.34)
$=\left(\amsmathbb{E}\Big{[}\exp\Big{(}\frac{is}{\sqrt{k}}\big{[}W^{\ast}-S_{t}(\varphi)\big{]}\Big{)}\Big{]}\right)^{k},$
where $W^{\ast}$ and $\{W_{i}\}_{i=1}^{k}$ were defined in the proof of Theorem 3.4. By (A.33), $\amsmathbb{E}[W^{\ast}-S_{t}(\varphi)]=0,$ therefore by another application of Lemma A.1, [34],
$\amsmathbb{E}\big{[}\exp(iz\,[W^{\ast}-S_{t}(\varphi)])\big{]}=1-\frac{z^{2}}{2}\amsmathbb{E}[(W^{\ast}-S_{t}(\varphi))^{2}]+\zeta(z,t),$ (A.35)
with $|\zeta(z,t)|\leq|z|^{2+\delta}\amsmathbb{E}|W^{\ast}-S_{t}(\varphi)|^{2+\delta}$ and some $\delta>0$ which can be chosen depending on the values of $\gamma_{G}$ and $\varepsilon$ according to Condition 1. As in the proof of Theorem 3.4, let us show that
$\sup_{t\geq T}\amsmathbb{E}|W^{\ast}-S_{t}(\varphi)|^{2+\delta}<\infty$ (A.36)
for some $T\geq 1$ and $\delta>0.$ Indeed, by Minkowski’s inequality,
$\left(\amsmathbb{E}|W^{\ast}-S_{t}(\varphi)|^{2+\delta}\right)^{\frac{1}{2+\delta}}\leq|S_{t}(\varphi)|+\left(\amsmathbb{E}|W^{\ast}|^{2+\delta}\right)^{\frac{1}{2+\delta}},$
where $|S_{t}(\varphi)|$ is bounded uniformly for large $t$ by slight modification of (A.33) and $\amsmathbb{E}\big{[}|W^{\ast}|^{2+\delta}\big{]}$ can be handled similarly to the proof of Theorem 3.4.
Therefore, returning to (A.34), we derive that for every $s\in\amsmathbb{R}$, as $k\to\infty$ uniformly in $t\geq T$,
$c_{k}^{t}(s)\to\exp(-s^{2}\amsmathbb{E}[(W^{\ast}-S_{t}(\varphi))^{2}]/2).$
By (A.36), the argument used for proving (A.8), and the dominated convergence theorem, we have
$\amsmathbb{E}[(W^{\ast}-S_{t}(\varphi))^{2}]\to\amsmathbb{E}[(W^{\circ}-S_{\circ}(\varphi))^{2}]={\rm var}(W^{\circ}),\quad t\to\infty.$
Repeating the last argument in the proof of Theorem 3.4 completes the proof. ∎
Proof of Theorem 3.7.
Let $\xi$ be a Pareto$(1)$-distributed random variable, independent of $\{Z_{i}\}_{i=1}^{n}.$ Without loss of generality, let us consider only $t$ such that $F$ is strictly monotone in the left neighborhood of $t.$ It is straightforward to verify that $F^{t}$ is the cdf of the random variable $U_{F}(s\xi)/U_{F}(s),$ where $s=U_{F}^{\leftarrow}(t)=(1-F(t))^{-1},$ see also (A.10). Indeed,
${\amsmathbb{P}}(U_{F}(s\xi)/U_{F}(s)\leq x)={\amsmathbb{P}}\big{(}\xi\leq s^{-1}U_{F}^{\leftarrow}(xt)\big{)}=1-\frac{1-F(xt)}{1-F(t)}=F^{t}(x).$ (A.37)
Denote
$\mu(s)=\amsmathbb{E}\,\varphi\left(\frac{U_{F}(s\xi)}{U_{F}(s)}\right).$
Thus, (3.13) can be rewritten as
$\sqrt{k}\big{(}\mu(U_{F}^{\leftarrow}(Z_{n-k,n}))-\mu(U_{F}^{\leftarrow}(U_{H}(n/k)))\big{)}\stackrel{{\scriptstyle P}}{{\to}}0,$
which is in accordance with relation (4.6) in [34]. Denote for brevity $u(s,\xi)=U_{F}(s\xi)/U_{F}(s).$ By the mean value theorem,
$|\mu(x)-\mu(y)|\leq\amsmathbb{E}\left[\left|\varphi^{\prime}(x^{\ast})\big{(}u(x,\xi)-u(y,\xi)\big{)}\right|\right],$ (A.38)
where $x^{\ast}$ lies between $u(x,\xi)$ and $u(y,\xi).$
Consider the asymptotic behaviour of the expression within the modulo on the right-hand side of the latter relation. Recall that a regularly varying function $U$ with index $\gamma$ is normalised if there exists a $T>1$ and a real measurable function $\gamma(\cdot)$ on $[T;\infty)$ converging to $\gamma$, such that for $t\geq T,$
$U(t)=U(T)\exp\left(\int_{T}^{t}\gamma(u)\frac{\mathop{}\!\mathrm{d}u}{u}\right).$ (A.39)
Clearly, the latter is satisfied for $U_{F}$ with $\gamma(t)=tU_{F}^{\prime}(t)/U_{F}(t)$ under the condition (3.10), and thus the conditions of Lemma 4.3 in [34] hold. Also note that we can choose an envelope $\bar{\varphi^{\prime}}$ such that $\bar{\varphi^{\prime}}(x)\neq 0,$ $x\geq 1.$ By the first assertion of the latter lemma and properties of the envelope $\bar{\varphi^{\prime}}$, we get
$\bar{\varphi^{\prime}}(x^{\ast})\leq C\max(\bar{\varphi^{\prime}}(\xi^{\gamma_{F}-\varepsilon^{\prime}}),\bar{\varphi^{\prime}}(\xi^{\gamma_{F}+\varepsilon^{\prime}}))$ (A.40)
for every $\varepsilon^{\prime}>0,$ some $C>0$ independent of $\varepsilon^{\prime},$ and large $x$ and $y$. Next, consider the difference within the modulo on the right-hand side of (A.38). By the relation (4.9) in Lemma 4.3, [34], we get for the above $\varepsilon^{\prime}$ and large $x$ and $y$
$\Big{|}u\big{(}x,\xi\big{)}-u\big{(}y,\xi\big{)}\Big{|}\leq 2\varepsilon^{\prime}|\log(x/y)|\xi^{\gamma_{F}+\varepsilon^{\prime}}.$ (A.41)
Let $\{\xi_{i}\}_{i=1}^{n}$ be independent Pareto$(1)$-distributed random variables, and $\xi_{1,n}\leq\ldots\leq\xi_{n,n}$ be the order statistics of this sample such that $Z_{n-k,n}=U_{H}(\xi_{n-k,n}).$ Since the function $U_{F}^{\leftarrow}(U_{H}(x))$ tends to infinity as $x\to\infty,$ we derive by (A.38), (A.40), and (A.41) for large $n$
$\sqrt{k}\Big{|}\mu(U_{F}^{\leftarrow}(Z_{n-k,n}))-\mu(U_{F}^{\leftarrow}(U_{H}(n/k)))\Big{|}$
$\leq 2\varepsilon^{\prime}C\sqrt{k}\left|\log\left(\frac{U_{F}^{\leftarrow}(U_{H}(\xi_{n-k,n}))}{U_{F}^{\leftarrow}(U_{H}(n/k))}\right)\right|\amsmathbb{E}\Big{[}\xi^{\gamma_{F}+\varepsilon^{\prime}}\max\big{(}\bar{\varphi^{\prime}}(\xi^{\gamma_{F}-\varepsilon^{\prime}}),\bar{\varphi^{\prime}}(\xi^{\gamma_{F}+\varepsilon^{\prime}})\big{)}\Big{]}.$
The mathematical expectation in the latter relation is bounded for small $\varepsilon^{\prime}$ by Condition 2. Thus, if we show that
$\sqrt{k}\log\left(\frac{U_{F}^{\leftarrow}(U_{H}(\xi_{n-k,n}))}{U_{F}^{\leftarrow}(U_{H}(n/k))}\right)=O_{\mathbb{P}}(1)$ (A.42)
then the proof is complete by letting $\varepsilon^{\prime}\to 0.$
But Smirnov’s lemma, [37], implies that $\xi_{n-k,n}$ is asymptotically normal as
$\frac{k^{3/2}}{n}(\xi_{n-k,n}-n/k)\stackrel{{\scriptstyle d}}{{\to}}N(0,1).$ (A.43)
Moreover, by (3.10) and (3.11), the derivative of the function $f(x)=\log(U_{F}^{\leftarrow}(U_{H}(x)))=-\log(1-F(U_{H}(x)))$ possesses the following property
$\lim_{x\to\infty}xf^{\prime}(x)=\lim_{x\to\infty}\frac{U_{H}(x)F^{\prime}(U_{H}(x))}{1-F(U_{H}(x))}\cdot\frac{xU_{H}^{\prime}(x)}{U_{H}(x)}=\frac{\gamma_{H}}{\gamma_{F}},$
therefore we derive by (A.43) and the delta-method,
$\sqrt{k}\,\frac{\gamma_{F}}{\gamma_{H}}\log\left(\frac{U_{F}^{\leftarrow}(U_{H}(\xi_{n-k,n}))}{U_{F}^{\leftarrow}(U_{H}(n/k))}\right)\stackrel{{\scriptstyle d}}{{\to}}N(0,1).$
Hence (A.42) follows, which completes the proof. ∎
Proof of Theorem 3.9.
The proof technique is inspired by Theorem 4.5 in [34]. Using the notation of the proof of Theorem 3.7, we get
$\sqrt{k}\int\varphi\mathop{}\!\mathrm{d}(F^{Z_{n-k,n}}-F^{\circ})=\sqrt{k}\big{(}\mu(U_{F}^{\leftarrow}(Z_{n-k,n}))-\mu(\infty)\big{)},$
where
$\mu(\infty):=\amsmathbb{E}\varphi(\xi^{\gamma_{F}})=\int\varphi\mathop{}\!\mathrm{d}F^{\circ}.$
Denoting $u(\infty,\xi)=\xi^{\gamma_{F}},$ we have by (A.38)
$\mu(x)-\mu(\infty)=\amsmathbb{E}\left[\varphi^{\prime}(x^{\ast})\big{(}u(x,\xi)-u(\infty,\xi)\big{)}\right],$ (A.44)
with $x^{\ast}$ lying between $u(x,\xi)$ and $u(\infty,\xi).$ By Potter’s bounds, for every $\varepsilon^{\prime}>0$ and large $t,$
$\frac{1}{2}x^{\gamma_{F}-\varepsilon^{\prime}}\leq\frac{U_{F}(tx)}{U_{F}(t)}\leq 2x^{\gamma_{F}+\varepsilon^{\prime}},$
thus similar to (A.40) and assuming w.l.o.g. that the envelope $\bar{\varphi^{\prime}}$ is positive
$\bar{\varphi^{\prime}}(x^{\ast})\leq C\max(\bar{\varphi^{\prime}}(0.5\,\xi^{\gamma_{F}-\varepsilon^{\prime}}),\bar{\varphi^{\prime}}(2\xi^{\gamma_{F}+\varepsilon^{\prime}})).$ (A.45)
for some $C>0.$ Next, the conditions of Lemma 4.6, [34], are satisfied under Condition 3, and therefore for some $K>0$ and large $x,$
$|u(x,\xi)-u(\infty,\xi)|=\left|\frac{U_{F}(x\xi)}{U_{F}(x)}-\xi^{\gamma_{F}}\right|\leq K\xi^{\gamma_{F}+\varepsilon^{\prime}}|a(x)|.$ (A.46)
Piecing together (A.44-A.46) yields
$\sqrt{k}\Big{|}\mu(U_{F}^{\leftarrow}(Z_{n-k,n}))-\mu(\infty)\Big{|}$
$\leq KC\sqrt{k}\,\big{|}a(U_{F}^{\leftarrow}(Z_{n-k,n}))\big{|}\,\amsmathbb{E}\Big{[}\xi^{\gamma_{F}+\varepsilon^{\prime}}\max\big{(}\bar{\varphi^{\prime}}(0.5\,\xi^{\gamma_{F}-\varepsilon^{\prime}}),\bar{\varphi^{\prime}}(2\xi^{\gamma_{F}+\varepsilon^{\prime}})\big{)}\Big{]},$ (A.47)
for large $n,$ where the mathematical expectation in the latter relation, similarly to the analogous step in the proof of Theorem 3.7, is bounded for small $\varepsilon^{\prime}$ by Condition 2. Since the functions $a,$ $U_{F}^{\leftarrow}=(1-F)^{-1}$ and $U_{H}$ are regularly varying and $\xi_{n-k,n}/(n/k)\stackrel{{\scriptstyle P}}{{\to}}1$ by Smirnov’s lemma (A.43), we have
$\frac{a(U_{F}^{\leftarrow}(Z_{n-k,n}))}{a(U_{F}^{\leftarrow}(U_{H}(n/k)))}=\frac{a(U_{F}^{\leftarrow}(U_{H}(\xi_{n-k,n})))}{a(U_{F}^{\leftarrow}(U_{H}(n/k)))}\stackrel{{\scriptstyle P}}{{\to}}1,$ (A.48)
which completes the proof in the case $\lambda=0.$
Along with this, for $\lambda\neq 0$ we get
$\varphi^{\prime}(x^{\ast})\stackrel{{\scriptstyle a.s.}}{{\to}}\varphi^{\prime}(\xi^{\gamma_{F}}),\quad\frac{u(x,\xi)-u(\infty,\xi)}{a(x)}\stackrel{{\scriptstyle a.s.}}{{\to}}\xi^{\gamma_{F}}h_{\rho}(\xi),\quad x\to\infty,$ (A.49)
where the former follows by the continuous mapping theorem and a.e. continuity of $\varphi^{\prime},$ and Condition 3 implies the latter. Therefore, by (A.45-A.47) and the dominated convergence theorem we derive
$\frac{\mu(x)-\mu(\infty)}{a(x)}\to\amsmathbb{E}\big{[}\varphi^{\prime}(\xi^{\gamma_{F}})\xi^{\gamma_{F}}h_{\rho}(\xi)\big{]}=C(\gamma_{F},\rho),\quad x\to\infty.$
Finally, by the latter convergence and (A.48)
$\sqrt{k}\big{(}\mu(U_{F}^{\leftarrow}(Z_{n-k,n}))-\mu(\infty)\big{)}\stackrel{{\scriptstyle P}}{{\to}}C(\gamma_{F},\rho)\lambda.$
The proof is complete. ∎
Appendix B Proofs of Section 4
Proof of Theorem 4.4.
The idea of the proof is similar to Theorem 2.1 and Theorem 4.1 in [34]. Let us first prove consistency of $R^{f}_{k,n}.$ It is known that the random variables $\{\xi_{n-j+1,n}/\xi_{n-k,n}\}_{j=1}^{k}$ are independent of $\xi_{n-k,n}$ and distributed like $\{\xi^{\ast}_{k-j+1,k}\}_{j=1}^{k},$ where $\xi_{1}^{\ast},\ldots,\xi_{k}^{\ast}$ are independent Pareto$(1)$-distributed. Thus, we have for the characteristic function of $R^{f}_{k,n}$
$\amsmathbb{E}\big{[}\exp(isR^{f}_{k,n})\big{]}=\amsmathbb{E}\left[\exp\left\{\frac{is}{k}\sum_{j=1}^{k}\vartheta\Big{(}f\Big{(}\frac{\xi_{n-j+1,n}}{\xi_{n-k,n}},\xi_{n-k,n}\Big{)}\Big{)}\right\}\right]=\amsmathbb{E}\big{[}a_{k}(\xi_{n-k,n})\big{]},$
where $s\in\amsmathbb{R}$ and
$a_{k}(t)=\amsmathbb{E}\left[\exp\left\{\frac{is}{k}\sum_{j=1}^{k}\vartheta\big{(}f\big{(}\xi_{j}^{\ast},t\big{)}\big{)}\right\}\right].$
Note that for fixed $t$, $\big{\{}\vartheta(f(\xi_{j}^{\ast},t))\big{\}}_{j=1}^{k}$ are independent random variables distributed like $\vartheta(f(\xi,t)),$ therefore
$a_{k}(t)=\left(\amsmathbb{E}\left[\exp\Big{\{}\frac{is}{k}\vartheta(f(\xi,t))\Big{\}}\right]\right)^{k}.$
Similarly to (A.31), we get for $z\in\amsmathbb{R}$
$\amsmathbb{E}\big{[}\exp(iz\,\vartheta(f(\xi,t)))\big{]}=1+iz\amsmathbb{E}\big{[}\vartheta(f(\xi,t))\big{]}+\eta(z,t),$
with $\eta(z,t)\leq 2|z|^{1+\delta}\amsmathbb{E}|\vartheta(f(\xi,t))|^{1+\delta}$ for $\delta>0$ from (4.6). By (4.4), continuity and eventual monotonicity of $\bar{\vartheta},$ assuming as before that $\bar{\vartheta}$ is positive, we derive for $t\geq T$ with $T$ from Definition 4
$|\vartheta(f(\xi,t))|\leq C\max\big{[}\bar{\vartheta}(g_{1}(\xi)),\bar{\vartheta}(g_{2}(\xi))\big{]},$ (B.1)
and thus by (4.6)
$\limsup_{t\to\infty}\amsmathbb{E}\big{[}|\vartheta(f(\xi,t))|^{1+\delta}\big{]}<\infty.$ (B.2)
Summarizing the above, we have
$a_{k}(t)\to\exp\big{\{}is\amsmathbb{E}[\vartheta(f(\xi,t))]\big{\}},\quad k\to\infty,$
uniformly in $t\geq T.$ Note that by (4.3) and a.e. continuity of $\vartheta$, $\vartheta(f(\xi,t))\stackrel{{\scriptstyle a.s.}}{{\to}}\vartheta(g(\xi)),t\to\infty,$ and therefore by (B.1), (4.6) and the dominated convergence theorem,
$\mu_{f}(t)=\amsmathbb{E}\big{[}\vartheta(f(\xi,t))\big{]}\to\amsmathbb{E}[\vartheta(g(\xi))],\quad t\to\infty.$ (B.3)
Appealing to the latter and repeating the last argument in the proof of Theorem 3.4 completes the proof of consistency.
Next, we prove (4.7). Let us consider the characteristic function of $\sqrt{k}\big{(}R^{f}_{k,n}-\mu_{f}(\xi_{n-k,n})\big{)}.$ We have for $s\in\amsmathbb{R}$
$\amsmathbb{E}\Big{[}\exp\big{(}is\sqrt{k}(R^{f}_{k,n}-\mu_{f}(\xi_{n-k,n}))\big{)}\Big{]}=\amsmathbb{E}\big{[}b_{k}(\xi_{n-k,n})\big{]},$
where
$b_{k}(t)=\amsmathbb{E}\left[\exp\left\{\frac{is}{\sqrt{k}}\sum_{j=1}^{k}\left[\vartheta\big{(}f\big{(}\xi_{j}^{\ast},t\big{)}\big{)}-\mu_{f}(t)\right]\right\}\right].$
Denote $Z(t)=\vartheta\big{(}f\big{(}\xi,t\big{)}\big{)}-\mu_{f}(t)$ and $Z_{j}^{\ast}(t)=\vartheta\big{(}f\big{(}\xi_{j}^{\ast},t\big{)}\big{)}-\mu_{f}(t),$ $j=1,\ldots,k.$ As above, for fixed $t$$\{Z_{j}(t)\}_{j=1}^{k}$ are independent zero-mean random variables distributed like $Z(t),$ and thus we can rewrite
$b_{k}(t)=\amsmathbb{E}\left[\exp\left\{\frac{is}{\sqrt{k}}\sum_{j=1}^{k}Z_{j}^{\ast}(t)\right\}\right]=\left(\amsmathbb{E}\left[\exp\left\{\frac{is}{\sqrt{k}}Z(t)\right\}\right]\right)^{k}.$ (B.4)
Similarly to (A.35), we get for $z\in\amsmathbb{R}$
$\amsmathbb{E}\big{[}\exp(iz\,Z(t))\big{]}=1-\frac{z^{2}}{2}\amsmathbb{E}[(Z(t))^{2}]+\zeta(z,t),$ (B.5)
with $|\zeta(z,t)|\leq|z|^{2+\delta^{\prime}}\,\amsmathbb{E}|Z(t)|^{2+\delta^{\prime}}$ for some $\delta^{\prime}>0$. By application of (4.4), (4.6) with $\delta^{\prime}=\delta-1$ and $\delta>1$ defined in Theorem 4.4, and (B.3), we get
$\limsup_{t\to\infty}\amsmathbb{E}\big{[}|Z(t)|^{2+\delta^{\prime}}\big{]}<\infty.$
Piecing together (B.4), (B.5) and the latter relation, we derive
$b_{k}(t)\to\exp\big{\{}-\amsmathbb{E}[Z^{2}(t)]s^{2}/2\big{\}},\quad k\to\infty,$
uniformly in $t\geq T.$ Along with this, by an argument similar to that we used to prove (B.3), we get by the dominated convergence theorem
$\amsmathbb{E}[Z^{2}(t)]\to{\rm var}\big{[}\vartheta(g(\xi))\big{]},\quad t\to\infty.$
Repeating the last argument in the proof of Theorem 3.4 completes the proof.
Proof of Theorem 4.5.
The proof follows along the lines of Theorem 3.9. Taking into account Theorem 4.4, it suffices to show that
$\sqrt{k}\big{(}\mu_{f}(\xi_{n-k,n})-\amsmathbb{E}[\vartheta(g(\xi))]\big{)}\stackrel{{\scriptstyle P}}{{\to}}C\,\lambda,$ (B.6)
Recall the notation $\mu_{f}(\infty):=\amsmathbb{E}[\vartheta(g(\xi))].$ Similarly to (A.38) and (A.44), we get
$\mu_{f}(t)-\mu_{f}(\infty)=\amsmathbb{E}\big{[}\vartheta^{\prime}(t^{\ast})\big{(}f(\xi,t)-g(\xi)\big{)}\big{]},$ (B.7)
where $t^{\ast}$ lies between $f(\xi,t)$ and $g(\xi).$ By (4.4) and assuming that $\bar{\vartheta^{\prime}}$ is positive as before, we get for $t\geq T$ by the properties of envelope
$\bar{\vartheta^{\prime}}(t^{\ast})\leq C\max\big{(}\bar{\vartheta^{\prime}}(g_{1}(\xi)),\bar{\vartheta^{\prime}}(g_{2}(\xi))\big{)}$
for some $C>0.$ Taking into account (4.9), we derive the following bound
$\sqrt{k}\left|\mu_{f}(\xi_{n-k,n})-\mu_{f}(\infty)\right|\leq C\sqrt{k}|\hat{a}(\xi_{n-k,n})|\,\amsmathbb{E}\Big{[}\hat{g}(\xi)\max\big{\{}\bar{\vartheta^{\prime}}(g_{1}(\xi)),\bar{\vartheta^{\prime}}(g_{2}(\xi))\big{\}}\Big{]}$ (B.8)
for large $n,$ where the latter expectation is bounded by (4.10), which completes the proof in the case $\lambda=0.$
Next, for $\lambda\neq 0,$ the a.e. continuity of $\vartheta^{\prime}$ together with the continuous mapping theorem and (4.8) imply that
$\vartheta^{\prime}(t^{\ast})\stackrel{{\scriptstyle a.s.}}{{\to}}\vartheta^{\prime}(g(\xi)),\quad\frac{f(\xi,t)-g(\xi)}{\hat{a}(t)}\stackrel{{\scriptstyle a.s.}}{{\to}}\tilde{g}(\xi),\quad t\to\infty.$
Therefore by (B.7, B.8) and the dominated convergence theorem
$\frac{\mu_{f}(t)-\mu_{f}(\infty)}{\hat{a}(t)}\to\amsmathbb{E}\big{[}\tilde{g}(\xi)\vartheta^{\prime}(g(\xi))\big{]}=C,\quad t\to\infty.$
Finally, the latter convergence and limit relation $\sqrt{k}\hat{a}(\xi_{n-k,n})\to\lambda$ imply (B.6). The proof is complete. ∎
Proof of Proposition 4.10.
Let’s start from the case $\gamma_{F}>0.$ However, the proof is similar in all three cases. We have
$1-\tilde{F}(x)=1-F(U_{H}(x))=x^{-\alpha_{F}}\ell_{\tilde{F}}(x)$
with $\ell_{\tilde{F}}(x)$ slowly varying at infinity. By the definition of $U_{F},$ the latter implies
$U_{H}(x)=U_{F}\big{(}x^{\alpha_{F}}(\ell_{\tilde{F}}(x))^{-1}\big{)},$ (B.9)
which holds independently of the value of $\gamma_{F}.$ Next, $U_{F}$ is regularly varying with index $\gamma_{F},$ thus by the properties of regularly varying functions, see, e.g., Theorem 1.5.7 (ii) in [8], $U_{H}$ is regularly varying with index $\gamma_{F}\alpha_{F}>0$, which completes the proof in case $\gamma_{F}>0.$
Now assume $\gamma_{F}<0.$ It is easily follows from the second part of Example 2 that if $\alpha_{F}>0$ then $\tau_{F}\leq\tau_{G}$ (to prove this, the condition $\gamma_{G}<0$ is not necessary). Thus, $\tau_{H}=\tau_{F}.$ By (B.9), we derive
$U_{H}(x)=\tau_{F}-f\big{(}x^{\alpha_{F}}(\ell_{\tilde{F}}(x))^{-1}\big{)},$
where $f$ is a regularly varying function with index $\gamma_{F}$ by the properties of $U_{F}.$ Applying again Theorem 1.5.7 (ii), [8], we get that $U_{H}=\tau_{F}-h(x),$ where $h(x)$ is a regularly varying function with index $\gamma_{F}\alpha_{F}<0,$ which completes the proof in case $\gamma_{F}<0.$
Finally, for the case $\gamma_{F}=0,$ note that the positivity of $\alpha_{F}$ implies $\tau_{F}=\tau_{H}$ independently of the value of $\gamma_{F}.$ Then, the proof can be carried out similarly to the first and second cases if $\tau_{F}=\infty$ and $\tau_{F}<\infty,$ respectively. ∎
Proof of Theorem 4.11.
The proof relies on the proofs of Theorem 3.1, Theorem 3.3, Theorem 3.4, and Theorem 3.6 under replacement of $\{Z_{i}\}$ with $\{\xi_{i}\},$$\{Z_{n-i,n}/Z_{n-k,n}\}$ with $\{f(\xi_{n-i,n}/\xi_{n-k,n},\xi_{n-k,n})\},$ and $F,$ $G,$ $H$ etc. with their tilde-versions. We also use $g_{1}$ and $g_{2}$ to bound the argument of $\vartheta$ in the analogue of (A.8) and similar situations instead of Potter’s bounds. Consequently, we restrict ourselves to indicating the differences.
First, Lemma A.1 also holds for pairs the $\{(\xi_{i}^{\ast},\delta_{i}^{\ast})\}.$ Indeed, since $H$ is eventually continuous, for large $n$ we get $\xi_{n-i+1,n}=(1-H(Z_{n-i+1,n}))^{-1},$ $i=1,\ldots,k.$ Hence we can choose
$\xi_{i}^{\ast}=\frac{1-H(t)}{1-H(tV_{i}^{\ast})},\quad i=1,\ldots,k,$
where the joint distribution of $\{\xi^{\ast}_{k-i+1,k}\}_{i=1}^{k}$ coincides with the joint distribution of $\{\xi_{n-i+1,n}/\xi_{n-k,n}\}_{i=1}^{k},$ and therefore pairs $\{(\xi_{i}^{\ast},\delta_{i}^{\ast})\}$ are i.i.d as well.
The second difference relates to the argument used in the proof of Theorem 3.1 to show boundedness of certain integrals and first applied to prove (A.8). Indeed, we cannot use that argument as it is, since for that we would need to impose some technical conditions on $f,$ $g_{1}$ and $g_{2}$ like monotonicity and continuity. Let us start with the case of zero $\alpha_{F}$. Consider as an example the version of the relation (A.8) and assume $\alpha_{F}=0$. By the monotone density theorem, see e.g. Theorem 1.7.2 in [8],
$\lim_{t\to\infty}\frac{t\tilde{F}^{\prime}(t)}{1-\tilde{F}(t)}=0.$
By Potter’s bounds, the index of regular variation of $\tilde{F}^{\prime}$ cannot be more than $-1$ since the lower Potter bound must be integrable. Next, by Karamata’s theorem, see Theorem 1.5.11 ibid., it cannot be less than $-1,$ since otherwise $\alpha_{F}<0.$ Thus applying Potter’s bounds once more, we get for some positive $\delta$ and large $t$
$(\tilde{F}^{t})^{\prime}(w)=\frac{\tilde{F}^{\prime}(wt)}{\tilde{F}^{\prime}(t)}\cdot\frac{t\tilde{F}^{\prime}(t)}{1-\tilde{F}(t)}\leq w^{-1+\delta},\quad w\geq 1.$ (B.10)
If $\alpha_{F}$ is positive then the regular variation property of $(\tilde{F}^{t})^{\prime}$ follows immediately from the monotone density theorem, therefore, applying Potter’s bounds, we get
$|(\tilde{F}^{t})^{\prime}(w)|\leq Aw^{-\alpha_{F}-1+\delta},\quad w\geq 1,$ (B.11)
for some $A,\delta>0$ and large $t.$ Note that
$|\tilde{\theta}^{t}|=\int_{1}^{\infty}\int_{1}^{w}|\vartheta^{t}_{f}(w)|\frac{1}{(1-\tilde{G}^{t}(v))}\tilde{G}^{t}(\mathop{}\!\mathrm{d}v)(\tilde{F}^{t})^{\prime}(w)\mathop{}\!\mathrm{d}w.$
By (B.10-B.11) and Potter’s bound applied to $1-\tilde{G}^{t},$ we derive for large $t$ and some $C_{1},C_{2},\delta^{\prime}>0$ similarly to (A.9)
$|\vartheta^{t}_{f}(w)(\tilde{F}^{t})^{\prime}(w)|\int_{1}^{w}\frac{1}{(1-\tilde{G}^{t}(v))}\tilde{G}^{t}(\mathop{}\!\mathrm{d}v)\leq C_{1}I(w\in[1,1+\delta^{\prime}])$
$\quad+\,C_{2}\big{|}\bar{\vartheta}(f(w,t))(\tilde{F}^{t})^{\prime}(w)\big{|}\log(w)I(w>1+\delta^{\prime})$
$\quad\leq C_{1}+A\,C_{2}\max\big{(}\bar{\vartheta}(g_{1}(w)),\bar{\vartheta}(g_{2}(w))\big{)}\log(w)\,w^{-\alpha_{F}-1+\delta}.$ (B.12)
The latter function is integrable by Condition 6, and so the dominated convergence theorem can be applied to prove the convergence of $\tilde{\theta}^{t}.$
Proof of Theorem 4.12.
By Theorem 4.11, it suffices to show
$\sqrt{k}\big{(}\mu_{f}(\xi_{n-k,n})-\mu_{f}(\infty)\big{)}\stackrel{{\scriptstyle P}}{{\to}}\mu.$ (B.13)
We have,
$\mu_{f}(\xi_{n-k,n})-\mu_{f}(\infty)$ $=$ $\int\vartheta(f(y,\xi_{n-k,n}))\tilde{F}^{\xi_{n-k,n}}(\mathop{}\!\mathrm{d}y)-\int\vartheta(g(y))\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}y)$
$=$ $\int\big{(}\vartheta(f(y,\xi_{n-k,n}))-\vartheta(g(y))\big{)}\tilde{F}^{\xi_{n-k,n}}(\mathop{}\!\mathrm{d}y)+\int\vartheta_{g}\mathop{}\!\mathrm{d}(\tilde{F}^{\xi_{n-k,n}}-\tilde{F}^{\circ})$
=:absent:=:= : $I_{1}+I_{2}.$
The proof of the relation
$\sqrt{k}I_{2}\stackrel{{\scriptstyle P}}{{\to}}\lambda\int x\vartheta_{g}^{\prime}(x)h_{\rho}(x^{\alpha_{F}})\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}x)$
actually coincides with the proof of Theorem 3.9 under replacement of $\varphi,$ $\gamma_{F},$ $F^{\circ},$ etc. with $\vartheta_{g},$ $1/\alpha_{F},$ $\tilde{F}^{\circ},$ etc., respectively. So, let us consider $I_{1}.$ Similarly to (B.7- B.8), we have by (B.11) for large $n$ and some $C,C_{1}>0$,
$\sqrt{k}|I_{1}|$ $=$ $\sqrt{k}\int|\vartheta(f(y,\xi_{n-k,n}))-\vartheta(g(y))|\tilde{F}^{\xi_{n-k,n}}(\mathop{}\!\mathrm{d}y)$
$\leq$ $C\sqrt{k}\int\max\big{(}\bar{\vartheta}^{\prime}(g_{1}(y)),\bar{\vartheta}^{\prime}(g_{2}(y))\big{)}|f(y,\xi_{n-k,n})-g(y)|\tilde{F}^{\xi_{n-k,n}}(\mathop{}\!\mathrm{d}y)$
$\leq$ $C_{1}\sqrt{k}|\hat{a}(\xi_{n-k,n})|\cdot\int\hat{g}(y)\max\big{(}\bar{\vartheta}^{\prime}(g_{1}(y)),\bar{\vartheta}^{\prime}(g_{2}(y))\big{)}y^{-\alpha_{F}-1+\varepsilon^{\prime}}\mathop{}\!\mathrm{d}y,$
where the last expression is finite by the relation $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\hat{\lambda}$ and Condition 7 and vanishes if $\hat{\lambda}=0,$ which completes the proof for this case. Moreover, (B) implies
$\frac{1}{|\hat{a}(t)|}\Big{|}\big{[}\vartheta(f(y,t))-\vartheta(g(y))\big{]}(\tilde{F}^{t})^{\prime}(y)\Big{|}\leq C_{1}\hat{g}(y)\max\big{(}\bar{\vartheta}^{\prime}(g_{1}(y)),\bar{\vartheta}^{\prime}(g_{2}(y))\big{)}y^{-\alpha_{F}-1+\varepsilon^{\prime}}.$ (B.15)
Next, for $\hat{\lambda}\neq 0,$ we get given $\xi_{n-k,n}=t$
$I_{1}=\int\big{(}\vartheta(f(y,t))-\vartheta(g(y))\big{)}(\tilde{F}^{t})^{\prime}(y)\mathop{}\!\mathrm{d}y$
$=\int\vartheta^{\prime}(t^{\ast})\big{(}f(y,t)-g(y)\big{)}(\tilde{F}^{t})^{\prime}(y)\mathop{}\!\mathrm{d}y,$
where $t^{\ast}$ lies between $f(y,t)$ and $g(y).$ The regular variation property of $(\tilde{F})^{\prime}$ and a.e. continuity of $\vartheta^{\prime}$ together with the continuous mapping theorem imply for every $y\geq 1$
$\vartheta^{\prime}(t^{\ast})\stackrel{{\scriptstyle a.e.}}{{\to}}\vartheta^{\prime}(g(y)),\quad(\tilde{F}^{t})^{\prime}(y)\to(\tilde{F}^{\circ})^{\prime}(y),\quad t\to\infty,$
where $\stackrel{{\scriptstyle a.e.}}{{\to}}$ denotes convergence almost everywhere. Therefore, by (B-B.15), the dominated convergence theorem and Condition 4,
$\frac{1}{\hat{a}(t)}I_{1}\to\int\vartheta^{\prime}(g(y))\tilde{g}(y)\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}y),\quad t\to\infty,$
which together with the limit relation $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\hat{\lambda}$ implies
$\sqrt{k}I_{1}\stackrel{{\scriptstyle P}}{{\to}}\hat{\lambda}\int\vartheta^{\prime}(g(y))\tilde{g}(y)\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}y).$
The proof is complete. ∎
Appendix C Proofs of Section 5
Proof of Theorem 5.1.
First, let us consider the case $\gamma_{F}>0.$ Then $\alpha_{F}$ is positive and we derive by Proposition 4.10 that $G$ and $H$ have regularly varying tails as well, and thus Theorem 3.4 can be applied. The assumption $\alpha_{F}>1/2$ (see also the remark below Theorem 5.1) implies Condition 1 for the functions $\varphi_{\ell}(x)=(\log x)^{\ell},$ $\ell=1,2,$ hence the assumptions of Theorem 3.4 hold for these two functions, obtaining
[…]
Thus, the result follows from the properties of convergence in probability and continuous mapping theorem.
The proof for the case $\gamma_{F}\leq 0$ is a bit more challenging and is based on the application of Theorem 4.11. Note that since $\alpha_{F}>0$ Proposition 4.10 implies $\gamma_{H}\leq 0$ and $x^{\ast}_{H}=x_{F}^{\ast}.$ As in Section 4.2, let us consider a candidate function
$f(x,t)=\frac{\log U_{H}(tx)-\log U_{H}(t)}{a_{H}(t)/U_{H}(t)},$ (C.1)
and therefore by (4.15) $g(x)=h_{\gamma_{H}}(x),$ where $h_{\gamma}$ is defined by (3.16). The candidates for $g_{1}$ and $g_{2}$ can be found in Section 4.2 as well, see (4.16) for the case $\gamma_{F}<0$ (one should replace $\gamma$ with $\gamma_{H}$ there), whereas for $\gamma_{F}=0$ one can use 0 and $\delta x^{\delta}$ as $g_{1}$ and $g_{2},$ respectively, for some $\delta\in(0,1)$ since $f(x,t)$ is increasing with respect to $x,$ see Remark 4.3.
Consider $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2}.$ By the relation $\alpha_{F}+\alpha_{G}=1$ we get $\alpha_{G}<1/2$ and thus, Condition 6 holds for $\vartheta_{1}$ and $\vartheta_{2},$ but one should select a properly small $\delta$ in case $\gamma_{F}=0.$ Next, the condition on existence of eventually monotone density for $\tilde{F}$ can be omitted because it is only needed to prove (B.12) and similar relations. Since $f(w,t)$ is monotone by $w$ and $\tilde{F}$ is regularly varying, the argument similar to the proof of (A.8) can be used to prove the convergence of $\tilde{\theta}^{t}$ and similar expressions. Therefore, applying Theorem 4.11 with necessary changes indicated above, we get
[…]
where $\xi$ is a Pareto(1) random variable and
$\amsmathbb{E}\big{(}h_{\gamma_{F}}(\xi)\big{)}=\frac{1}{1-\gamma_{F}},\quad\amsmathbb{E}\big{(}h_{\gamma_{F}}(\xi)\big{)}^{2}=\frac{2}{(1-\gamma_{F})(1-2\gamma_{F})}.$
An application of the properties of convergence in probability and the continuous mapping theorem completes the proof. ∎
Proof of Theorem 5.2.
First, let us prove (i). We start with the following key two-dimensional result.
Lemma C.1.
Assume the conditions of Theorem 5.2 (i). Then
[…]
where
$b=\left(\begin{array}[]{c}\frac{1}{1-\rho}\\ \frac{2\gamma_{F}(2-\rho)}{(1-\rho)^{2}}\end{array}\right),\quad\Sigma=\left(\begin{array}[]{cc}\frac{\gamma_{G}\gamma_{F}^{2}}{\gamma_{G}-\gamma_{F}}&\frac{2\gamma_{G}\gamma_{F}^{3}(2\gamma_{G}-\gamma_{F})}{(\gamma_{G}-\gamma_{F})^{2}}\\ \frac{2\gamma_{G}\gamma_{F}^{3}(2\gamma_{G}-\gamma_{F})}{(\gamma_{G}-\gamma_{F})^{2}}&\frac{4\gamma_{G}\gamma_{F}^{4}(5\gamma_{G}^{2}-4\gamma_{G}\gamma_{F}+\gamma_{F}^{2})}{(\gamma_{G}-\gamma_{F})^{3}}\end{array}\right).$
Proof.
To prove the lemma, we require a two-dimensional analogue of Theorem 3.6, since then all other results of Section 3 concern convergence in probability and hence they are extended to the two-dimensional case straightforwardly. Thus, assume two functions $\varphi^{(1)}$ and $\varphi^{(2)}$ satisfying the assumptions of Theorem 3.6. Conducting the proof of Theorem 3.1 and Theorem 3.3 for the pair $\big{(}S_{k,n}(\varphi^{(1)}),S_{k,n}(\varphi^{(2)})\big{)}$ together, we get similarly to (A.29) the following decomposition:
$\big{(}S_{k,n}(\varphi^{(1)}),\,S_{k,n}(\varphi^{(2)})\big{)}=\left(\frac{1}{k}\sum_{j=1}^{k}W_{j}^{(1)}+r^{(1)}_{k,n},\;\frac{1}{k}\sum_{j=1}^{k}W_{j}^{(2)}+r^{(2)}_{k,n}\right),$
where for every $j\in\{1,\ldots,k\}$ $W_{j}^{(1)}$ and $W_{j}^{(2)}$ given $Z_{n-k,n}=t$ are the functions of $V^{\ast}_{j}$ and $\delta_{j}^{\ast}$ defined similarly to (A.30) by replacing $V^{\ast},\delta^{\ast}$ with $V^{\ast}_{j},\delta^{\ast}_{j}$ and $\varphi$ with $\varphi^{(1)}$ and $\varphi^{(2)},$ respectively. So $\big{\{}(W_{j}^{(1)},W_{j}^{(2)})\big{\}}_{j=1}^{k}$ are i.i.d. given $Z_{n-k,n}=t,$ and both $\sqrt{k}r^{(1)}_{k,n},$ $\sqrt{k}r^{(2)}_{k,n}$ are $\overline{o}_{P}(1)$ and thus do not affect the asymptotic. Following the proof of Theorem 3.6 in two dimensions (in fact, the only required change is to write $\langle\vec{s},\vec{\varphi}\rangle=s^{(1)}\varphi^{(1)}+s^{(2)}\varphi^{(2)}$ instead of $s\varphi$), we get that the vector
$\sqrt{k}\Big{(}S_{k,n}(\varphi^{(1)})-S_{Z_{n-k,n}}(\varphi^{(1)}),\,S_{k,n}(\varphi^{(2)})-S_{Z_{n-k,n}}(\varphi^{(2)})\Big{)}$
converges in distribution to the zero-mean Gaussian random vector with covariance matrix $\Sigma$ given as the covariance matrix of the vector $\big{(}W^{\circ}(\varphi^{(1)}),W^{\circ}(\varphi^{(2)})\big{)}$ built according to formula (3.8) and using the same $V^{\circ}$ and $\delta^{\circ}.$
The result of the lemma now follows by a routine calculation of the asymptotic bias (following from Theorem 3.9) and asymptotic variance for $\varphi^{(1)}(x)=\log x$ and $\varphi^{(2)}(x)=\log^{2}x.$
By an application of the above lemma, we obtain (i), appealing to Cramér’s delta method.
The result of (ii) follows immediately from Theorem 4.12, the analogue of Lemma C.1 for $\gamma_{F}=0$ (the proof technique of the two results is virtually identical), and Cramér’s delta method. The only detail worth clarifying is the calculation of the bias in Theorem 4.12. The function $\tilde{g}$ appearing in the bias term in this theorem is equal to $H_{\gamma_{-},\rho^{\prime}}(\cdot)$ by (4.17), where $\gamma_{-}=\rho^{\prime}=0$ by Lemma B.3.16 in [15]. Thus, $\tilde{g}(x)=\frac{1}{2}\log^{2}x$ by (B.3.8) ibid. The remaining part of the proof is straightforward.
Now, let us prove (iii). To obtain minimal conditions, we prove (4.25) once again for particular $f$ and $g.$
Since under the assumptions of Theorem 5.1, Theorem 4.11 holds for $f$ defined by (C.1) and $g=h_{\gamma_{H}}$, it is enough to show (B.13) in our particular case, see the corresponding steps in the proofs of Theorem 4.11 and Theorem 4.12. Denoting for the moment $s=\xi_{n-k,n},$ we have,
$\mu_{f}(s)-\mu_{f}(\infty)$ $=$ $\int_{1}^{\infty}\vartheta(f(y,s))\tilde{F}^{s}(\mathop{}\!\mathrm{d}y)-\int_{1}^{\infty}\vartheta(g(y))\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}y)$ (C.3)
$=$ $\int_{1}^{\infty}\vartheta\left(\frac{\log U_{H}(sy)-\log U_{H}(s)}{a_{H}(s)/U_{H}(s)}\right)\mathop{}\!\mathrm{d}\left(\frac{F(U_{H}(sy))}{1-F(U_{H}(s))}\right)$
$-\int_{1}^{\infty}\vartheta\left(\frac{y^{\gamma_{H}}-1}{\gamma_{H}}\right)\mathop{}\!\mathrm{d}(1-y^{-\alpha_{F}}),$
where $\alpha_{F}=\gamma_{H}/\gamma_{F}$ by Example 2 and Proposition 4.10. Denote for brevity $q:=a_{H}/U_{H}$ and recall that $\tau:=\tau_{F}=\tau_{H}$ by Proposition 4.10. By Theorem B.2.18 in [15], $q$ can be chosen equal to $-\gamma_{H}(\tau_{F}-U_{H}).$ Let us change the argument of the integral (C.3) to
$x=\frac{\tau-U_{H}(sy)}{\tau-U_{H}(s)},$
and the argument of the integral (C.3) to $x=y^{\gamma_{H}}.$ Notice that these changes are natural since the latter ratio tends to $y^{\gamma_{H}}$ by the properties of distributions belonging to the Weibull max-domain of attraction. Then, we can rewrite
$\mu_{f}(s)-\mu_{f}(\infty)$ $=$ $-\int_{0}^{1}\vartheta\left(\frac{\log(\tau-xt)-\log(\tau-t)}{-\gamma_{H}(\log\tau-\log(\tau-t))}\right)\mathop{}\!\mathrm{d}\left(\frac{F(\tau-xt)}{1-F(\tau-t)}\right)$
$+\int_{0}^{1}\vartheta\Big{(}\frac{x-1}{\gamma_{H}}\Big{)}\mathop{}\!\mathrm{d}(1-x^{-1/\gamma_{F}}),$
where $t:=\tau-U_{H}(s)$ tends to 00 as $s\to\infty.$ Denote $F^{\circ}(1-x)=1-x^{-1/\gamma_{F}},$
$F^{\ast t}(1-x)=\frac{F(\tau-xt)-F(\tau-t)}{1-F(\tau-t)}$
$\chi(x,t)=\frac{1}{\gamma_{H}}\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}-1\right),\quad\chi(x,0)=\frac{x-1}{\gamma_{H}}.$
Clearly, $\chi(x,0)=\lim_{t\downarrow 0}\chi(x,t).$ Using this notation, we get
$-(\mu_{f}(s)-\mu_{f}(\infty))$ $=$ $\int_{0}^{1}\vartheta(\chi(x,t))\mathop{}\!\mathrm{d}(F^{\ast t}(1-x))-\int_{0}^{1}\vartheta(\chi(x,0))\mathop{}\!\mathrm{d}(F^{\circ}(1-x))$
$=$ $\left(\int_{0}^{1}\vartheta(\chi(x,t))\mathop{}\!\mathrm{d}(F^{\ast t}(1-x))-\int_{0}^{1}\vartheta(\chi(x,0))\mathop{}\!\mathrm{d}(F^{\ast t}(1-x))\right)$
$+\left(\int_{0}^{1}\vartheta(\chi(x,0))\mathop{}\!\mathrm{d}[F^{\ast t}(1-x)-F^{\circ}(1-x)]\right)=:I_{1}+I_{2}.$
First, consider the asymptotic behaviour of $I_{2}.$ Similar to (A.37), notice that
$U_{F^{\ast t}}(x)=1-\frac{\tau-U_{F}(xv)}{\tau-U_{F}(v)},$
where $t=\tau-U_{F}(v),$ is the tail quantile function of $F^{\ast t}$ (here we think about $t$ and $v$ as constants) and that it converges to $1-x^{\gamma_{F}}$ as $v\to\infty$ by the regular variation property of $U_{F}.$ Hence $U_{F^{\ast t}}(\xi)$ has a cdf $F^{\ast t},$ where $\xi$ is a a Pareto(1) random variable, and
$v=U_{F}^{\leftarrow}(\tau-t)=U_{F}^{\leftarrow}(U_{H}(\xi_{n-k,n}))=U_{F}^{\leftarrow}(Z_{n-k,n})$
as in the proof of Theorem 3.9. The remaining part of finding the asymptotic behaviour of $I_{2}$ is a repetition of the steps of that proof (we let the analogue of Condition 2 hold) for $\varphi(x)=\vartheta(\chi(x,0)),$ $u(v,\xi)=1-U_{F^{\ast t}}(\xi),$ and replacement of Condition 3 by the second order condition (5.5). Therefore, we derive
[…]
where
$C_{2}(\gamma_{F},\rho)=\amsmathbb{E}\big{[}\varphi^{\prime}(\xi^{\gamma_{F}})\xi^{\gamma_{F}}h_{\rho}(\xi)\big{]}=\frac{1}{\gamma_{H}}\int_{1}^{\infty}\vartheta^{\prime}(\chi(x^{\gamma_{F}},0))x^{\gamma_{F}-2}h_{\rho}(x)\mathop{}\!\mathrm{d}x.$
Now, let us prove that
[…]
with
$C_{1}(\gamma_{F})=\int_{0}^{1}\vartheta^{\prime}\Big{(}\frac{x-1}{\gamma_{H}}\Big{)}\frac{x(x-1)}{2\gamma_{H}}\mathop{}\!\mathrm{d}(F^{\circ}(1-x)).$
Denote $\vartheta_{H}(x)=\vartheta((x-1)/\gamma_{H})$ for brevity. Keeping the notation $t=\tau-Z_{n-k,n},$ we have
$I_{1}$ $=$ $\int_{0}^{1}\left[\vartheta_{H}\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}\right)-\vartheta_{H}(x)\right]\mathop{}\!\mathrm{d}(F^{\ast t}(1-x)).$
Let us analyse the difference within the squared brackets in the latter integral. By Taylor expansion and the mean value theorem, we get uniformly in $x$
$\vartheta_{H}\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}\right)-\vartheta_{H}(x)$ $=$ $\vartheta_{H}^{\prime}(x^{\ast})\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}-x\right)$ (C.5)
$=$ $\vartheta_{H}^{\prime}(x^{\ast})\frac{x(x-1)}{2\tau}t(1+o_{P}(1)),$
where
$x^{\ast}\in\left[\frac{\log(1-xt/\tau)}{\log(1-t/\tau)},x\right],$
since the latter ratio of logarithms is less than $x$ for $x\in[0,1]$ and $t\in(0,\tau),$ which holds with probability tending to $1$ as $n\to\infty.$ Note also that for $x\in[0,1]$
$0<\frac{1}{1-xt/\tau}\leq C(1+o_{P}(1)),\quad n\to\infty,$
for some positive $C.$ On the other hand, using the simple bounds $\frac{y}{1+y}<\log(1+y)<y$ for $y>-1$ and assuming $|\vartheta_{H}^{\prime}|$ increases (as are $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2}$), we get
$\sqrt{k}|I_{1}|$$\leq$ $\sqrt{k}\int_{0}^{1}|\vartheta_{H}^{\prime}(x)|\left|\frac{x}{1-xt/\tau}-x\right|\mathop{}\!\mathrm{d}(F^{\ast t}(1-x))$ (C.6)
$\leq$ $C\sqrt{k}(1-Z_{n-k,n}/\tau)\int_{0}^{1}|\vartheta_{H}^{\prime}(x)|\,x^{2}\,x^{-1/\gamma_{F}-1+\varepsilon}\mathop{}\!\mathrm{d}x(1+o_{P}(1)),$
where the latter inequality holds by the argument used for proving (A.8). If
$\int_{0}^{1}|\vartheta_{H}^{\prime}(x)|\,x^{1-1/\gamma_{F}+\varepsilon}\mathop{}\!\mathrm{d}x<\infty,$ (C.7)
which holds for $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2},$ then the expression in the right-hand side of (C.6) is finite by the relation $\sqrt{k}(\tau-U_{H}(n/k))\to\hat{\lambda}\tau$ and the argument similar to (A.48), and, moreover, vanishes if $\hat{\lambda}=0,$ which completes the proof of (C.4) for this case.
Next, denote
$f_{1}(x,t)=\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}-x\right)\vartheta_{H}^{\prime}\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}\right),$
$f_{2}(x,t)=\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}-x\right)\vartheta_{H}^{\prime}(x),$
and assume $\vartheta_{H}^{\prime}$ is a.e. continuous (as is for $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2}$). Notice that
$\frac{1}{t}|f_{1}(x,t)|\leq\frac{1}{t}|f_{2}(x,t)|\leq Cx^{2}\vartheta_{H}^{\prime}(x)$
for small $t$; in the sequel we use these relations when applying the dominated convergence theorem. Given $\tau-Z_{n-k,n}=t,$ $\sqrt{k}I_{1}$ lies between $\sqrt{k}\amsmathbb{E}\big{[}f_{1}(1-U_{F^{\ast t}}(\xi),t)\big{]}$ and $\sqrt{k}\amsmathbb{E}\big{[}f_{2}(1-U_{F^{\ast t}}(\xi),t)\big{]},$ which both converge to $\hat{\lambda}C_{1}(\gamma_{F})$ as $t\downarrow 0.$ Indeed, by the dominated convergence theorem, (C.5-C.6), convergence $U_{F^{\ast t}}(x)\to U_{F^{\circ}}(x)$ and a.e. continuity of $f_{1}(x,t)$ in the first argument
$\frac{\tau}{t}\amsmathbb{E}\big{[}f_{1}(1-U_{F^{\ast t}}(\xi),t)\big{]}\to\int_{0}^{1}\vartheta^{\prime}_{H}(x)\frac{x(x-1)}{2}\mathop{}\!\mathrm{d}(F^{\circ}(1-x))=C_{1}(\gamma_{F}),\quad t\downarrow 0,$
and the same holds for $f_{2}.$ The above together with $\sqrt{k}(1-U_{H}(n/k)/\tau)\to\hat{\lambda}$ and an argument similar to (A.48) gives us (C.4).
Thus, for $\vartheta_{\ell}(x)=x^{\ell},$ $\ell=1,2,$ we get the following result
[…]
where
$b_{\ell}=\lambda\frac{\ell\gamma_{F}^{\ell-1}(1+(\ell-1)(1-\rho-3\gamma_{F}))}{\gamma_{H}^{\ell}\prod_{i=1}^{\ell}(1-i\gamma_{F})(1-\rho-i\gamma_{F})}-\hat{\lambda}\frac{\ell^{2}\gamma^{\ell}_{F}}{2\gamma_{H}^{\ell}\prod_{i=1}^{\ell+1}(1-i\gamma_{F})},$
$\sigma^{2}_{1}=\frac{(\gamma_{F}+\gamma_{G})^{2}}{\gamma_{G}(1-\gamma_{F})^{2}(\gamma_{G}(1-2\gamma_{F})-\gamma_{F})},$
$\sigma_{2}^{2}=\frac{4(\gamma_{F}+\gamma_{G})^{4}(22\gamma_{F}^{2}\gamma_{G}^{2}-21\gamma_{F}\gamma_{G}^{2}+5\gamma^{2}_{G}+9\gamma_{F}^{2}\gamma_{G}-4\gamma_{F}\gamma_{G}+\gamma^{2}_{F})}{\gamma_{G}^{3}(1-\gamma_{F})^{2}(1-2\gamma_{F})^{2}\prod_{j=2}^{4}(\gamma_{G}-\gamma_{F}-j\gamma_{F}\gamma_{G})}.$
Next, repeating the proof of Lemma C.1 with necessary changes, we get the two-dimensional version of (C.8) with the limit distribution $N(\vec{b},\Sigma),$ where $\vec{b}=(b_{1},b_{2})^{T},$ $\Sigma=\|\Sigma_{ij}\|_{i,j=1}^{2},$ $\Sigma_{11}=\sigma_{1}^{2},$ $\Sigma_{22}=\sigma_{2}^{2},$ and
$\Sigma_{12}=\frac{2(\gamma_{F}+\gamma_{G})^{3}(2\gamma_{G}-\gamma_{F}-4\gamma_{F}\gamma_{G})}{\gamma_{G}^{2}(1-\gamma_{F})^{2}(1-2\gamma_{F})\prod_{j=2}^{3}(\gamma_{G}-\gamma_{F}-j\gamma_{F}\gamma_{G})}.$
An appeal to Cramér’s delta method now yields (iii) with
$\sigma^{2}=a_{1}^{2}\sigma^{2}_{1}+2a_{1}a_{2}\Sigma_{12}+a^{2}_{2}\sigma^{2}_{2}.$ (C.9)
and
$a_{1}=1-\frac{2\gamma_{G}(1-\gamma_{F})^{2}(1-2\gamma_{F})}{\gamma_{F}+\gamma_{G}},\quad a_{2}=\frac{\gamma_{G}^{2}(1-\gamma_{F})^{2}(1-2\gamma_{F})^{2}}{2(\gamma_{F}+\gamma_{G})^{2}}.$